### v1

# Spatial-Sign based Maxsum Test for High Dimensional Location Parameters

Abstract
In this study, we explore a robust testing procedure for the high-dimensional location parameters testing problem. Initially, we introduce a spatial-sign based max-type test statistic, which exhibits excellent performance for sparse alternatives. Subsequently, we demonstrate the asymptotic independence between this max-type test statistic and the spatial-sign based sum-type test statistic (Feng and Sun, 2016). Building on this, we propose a spatial-sign based max-sum type testing procedure, which shows remarkable performance under varying signal sparsity. Our simulation studies underscore the superior performance of the procedures we propose.
Keywords: Asymptotic independence, High dimensional data, Scalar-invariant, Spatial-sign.
1 Introduction
The testing of location parameters is a crucial and extensively researched area in multivariate statistics with a fixed dimension. The conventional Hotelling’s $T^{2}$ test is commonly applied, but it fails in high-dimensional scenarios where the variable’s dimension $p$ exceeds the sample sizes $n$. Consequently, numerous efforts have been made to develop a high-dimensional mean test procedure. One straightforward approach is to substitute the Mahalanobis distance with the Euclidean distance. For the two-sample location problem, Bai and Saranadasa (1996) employed the $L_{2}$-norm of the difference between two sample means. Chen and Qin (2010) eliminated some redundant terms in Bai and Saranadasa (1996)’s test statistics and made no assumptions about the relationship between the dimension and sample sizes. Srivastava (2009), Park and Ayyala (2013), and Feng et al. (2015) suggested some scalar-invariant test statistics that replace the sample covariance matrix in Hotelling’s $T^{2}$ test statistics with its diagonal matrix. All these methods are built on the assumption of normal distribution or diverging factor models, which perform poorly for heavy-tailed distributions. For instance, the well-known multivariate t-distribution does not meet the above assumption. Therefore, numerous studies have also considered robust high-dimensional test procedures.
In traditional multivariate analysis, numerous methods have been developed to extend classic univariate rank and signed rank techniques to a multivariate context. A significant method is based on spatial signs and ranks, utilizing the so-called Oja median (Oja, 2010). Wang et al. (2015) proposed a high-dimensional spatial-sign test that replaces the scatter matrix with the identity matrix for a one-sample location problem. Similarly, Feng and Sun (2016) proposed a high-dimensional spatial sign test that replaces the scatter matrix with its diagonal matrix, which has a scalar-invariant property. Furthermore, Feng et al. (2016) considered the high-dimensional two-sample location problem based on the spatial-sign method. Feng, Liu, and Ma (2021) devised an inverse norm sign test that considers not only the direction of the observations but also the modulus of the observation. Huang et al. (2023) extended the inverse norm sign test for a high-dimensional two-sample location problem. Feng et al. (2020) demonstrated that the spatial-rank method also performs well for a high-dimensional two-sample problem. All these methods are constructed using the $L_{2}$-norm of the spatial median, which performs well under dense alternatives, meaning many variables have non-zero means. However, it is well-known that these sum-type test procedures perform poorly for sparse alternatives, where only a few variables have non-zero means.
In high-dimensional settings, numerous max-type test procedures have been introduced to detect sparse alternatives. Cai et al. (2013) proposed a test statistic for the high-dimensional two-sample mean problem, which is based on the maximum difference between the means of two samples’ variables under the Gaussian or sub-Gaussian assumption. For heavy-tailed distributions, Cheng et al. (2023) established a Gaussian approximation for the sample spatial median over the class of hyperrectangles and constructed a max-type test procedure using a multiplier bootstrap algorithm. However, their proposed test statistic is not scalar-invariant, and they did not provide the limit null distribution of their test statistic. The multiple bootstrap algorithm is also time-consuming. In this paper, we first introduce a novel Bahadur representation of the scaled sample spatial median and then construct a new max-type test statistic. We demonstrate that the limit null distribution of the proposed test statistic is still a Type I Gumbel distribution. We also establish the consistency of the proposed max-type test procedure. Simulation studies further illustrate its superiority over existing methods under sparse alternatives and heavy-tailed distributions.
In practical scenarios, it’s often unknown whether the alternative is dense or sparse. This has led to numerous studies proposing an adaptive strategy that combines the sum-type test and max-type test. For high-dimensional mean problems, Xu et al. (2016) integrated different $L_{r}$-norms of the sample means. He et al. (2021) introduced a family of U-statistics as an unbiased estimator of the $L_{r}$-norm of the mean vectors, demonstrating that U-statistics of different finite orders are asymptotically independent, normally distributed, and also independent from the maximum-type test statistic. Feng et al. (2022a) relaxed the covariance matrix assumption to establish independence between the sum-type test statistic and the max-type test statistic. There are also many other studies showing the asymptotic independence between the sum-type test statistics and the max-type test statistic for other high-dimensional problems. For instance, Wu et al. (2019) and Wu et al. (2020) examined the coefficient test in high-dimensional generalized linear models. Feng et al. (2022b) looked at the cross-sectional independence test in high-dimensional panel data models. Yu et al. (2022) focused on testing the high-dimensional covariance matrix. Feng et al. (2022d) considered the high-dimensional white noise test, while Wang and Feng (2023) looked at high-dimensional change point inference. Ma et al. (2023) and Yu et al. (2023) considered testing the alpha of high-dimensional time-varying and linear factor pricing models, respectively.
However, all these methods assume a normal or other light-tailed distributions. There’s a gap in the literature when it comes to considering the asymptotic independence between the sum-type test statistic and the max-type test statistic under heavy-tailed distributions. In this paper, we first establish the asymptotic independence between Feng and Sun (2016)’s sum-type test statistic and a newly proposed spatial sign-based max-type test statistic for high dimensional one sample location parameter problem. We then propose a Cauchy combination test procedure (Liu and Xie, 2020) to handle general alternatives. Both simulation studies and theoretical results demonstrate the advantages of our newly proposed methods.
This paper are organized as follow. Section 2 introduce Bahadur representation of the scaled spatial median and establish the max-type test statistic. In section 3, we prove the asymptotic independence between the sum-type test statistic and the new proposed max-type test statistic and construct the Cauchy combination test procedure. Section 4 show some simulation studies. We give a real data application in Section 5. Some discussion are stated in Section 6. All the technical details are in the Appendix.
Notations: For $d$-dimensional $\bm{x}$, we use the notation $\|\bm{x}\|$ and $\|\bm{x}\|_{\infty}$ to denote its Euclidean norm and maximum-norm respectively. Denote $a_{n}\lesssim b_{n}$ if there exists constant $C$, $a_{n}\leq Cb_{n}$ and $a_{n}\asymp b_{n}$ if both $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$ hold. Let $\psi_{\alpha}(x)=\exp\left(x^{\alpha}\right)-1$ be a function defined on $[0,\infty)$ for $\alpha>0$. Then the Orlicz norm $\|\cdot\|_{\psi_{\alpha}}$ of a $\bm{X}$ is defined as $\|\bm{X}\|_{\psi_{\alpha}}=\inf\left\{t>0,\mathbb{E}\left\{\psi_{\alpha}(|\bm{X}|/t)\right\}\leqslant 1\right\}$. Let $\operatorname{tr}(\cdot)$ be a trace for matrix, $\lambda_{min}(\cdot)$ and $\lambda_{max}(\cdot)$ be the minimum and maximum eigenvalue for symmetric martix. $\mathbf{I}_{p}$ represents a p-dimensional identity matrix, and $\operatorname{diag}\{v_{1},v_{2},\cdots,v_{p}\}$ represents the diagonal matrix with entries $\bm{v}=(v_{1},v_{2},\cdots,v_{p})$. For $a,b\in\mathbb{R}$, we write $a\wedge b=\min\{a,b\}$.
2 Max-type test
Let $\bm{X}_{1},\ldots,\bm{X}_{n}$ be a sequence of independent and identically distributed (i.i.d.) $p$-dimensional random vectors from a population $X$ with cumulative distribution function $F_{X}$ in $\mathbb{R}^{p}$. We consider the following model:
$\bm{X}_{i}=\bm{\theta}+v_{i}\mathbf{\Gamma}\bm{W}_{i},$ (2.1)
where $\bm{\theta}$ is location parameter, $\bm{W}_{i}$ is a p-dimensional random vector with independent components, $E(\bm{W}_{i})=0$, $\mathbf{\Sigma}=\mathbf{\Gamma}\mathbf{\Gamma}^{\top}$, $v_{i}$ is a nonnegative univariate random variable and is independent with the spatial sign of $\bm{W}_{i}$. The distribution of $\bm{X}$ depends on $\mathbf{\Gamma}$ through the shape matrix. Model (2.1) encompasses a wide range of frequently utilized multivariate models and distribution families, such as the independent components model (Nordhausen et al., 2009; Ilmonen and Paindaveine, 2011; Yao et al., 2015) and the family of elliptical distributions (Hallin and Paindaveine, 2006; Oja, 2010; Fang, 2018).
In this paper, we focus on the following one sample testing problem
$H_{0}:\bm{\theta}=\mathbf{0}\text{ versus }\quad H_{1}:\bm{\theta}\neq\mathbf{0}.$ (2.2)
The spatial sign function is defined as $U(\mathbf{x})=\|\mathbf{x}\|^{-1}\mathbf{x}I(\mathbf{x}\neq\mathbf{0})$. In traditional fixed $p$ circumstance, the following so-called ”inner centering and inner standardization” sign-based procedure is usually used (cf., Chapter 6 of Oja (2010))
$Q_{n}^{2}=np\overline{\bm{U}}^{T}\overline{\bm{U}},$
where $\overline{\bm{U}}=\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{U}}_{i},\hat{\bm{U}}_{i}=U\left(\mathbf{S}^{-1/2}\bm{X}_{i}\right),\mathbf{S}^{-1/2}$ are Tyler’s scatter matrix (cf., Section 6.1.3 of Oja (2010)). $Q_{n}^{2}$ is affine-invariant and can be regarded as a nonparametric counterpart of Hotelling’s $T^{2}$ test statistic by using the spatial-signs instead of the original observations $\bm{X}_{i}$ ’s. However, when $p>n,Q_{n}^{2}$ is not defined as the matrix $\mathbf{S}^{-1/2}$ is not available in high-dimensional settings.
In high-dimensional settings, Wang et al. (2015) proposed a method where Tyler’s scatter matrix is replaced by the identity matrix. This led to the following test statistic:
$T_{WPL}=\sum_{i<j}U^{T}(\bm{X}_{i})U(\bm{X}_{j}).$
Building on this, Feng and Sun (2016) extended the method and introduced a scalar-invariant spatial-sign based test procedure, which will be detailed in subsection 3.1. Both methods utilize sum-type test statistics, which perform well under dense alternatives where many elements of $\bm{\theta}$ are nonzeros. However, their power decreases under sparse alternatives where only a few elements of $\bm{\theta}$ are nonzeros.
It is well-known that max-type tests have good performance under sparse alternatives (Cai et al., 2013). Therefore, Cheng et al. (2023) first provided the Bahadur representation of the classic spatial median $\widetilde{\bm{\theta}}$, defined as:
$\widetilde{\bm{\theta}}=\arg\min_{\bm{\theta}}\sum_{i=1}^{n}||{\bf X}_{i}-\bm{\theta}||.$ (2.3)
They then proposed a max-type test procedure based on Gaussian approximation. While this approach is robust and effective in high-dimensional settings, it loses scalar information of different variables and is not scalar-invariant. In real-world scenarios, different components may have entirely different physical or biological readings, and their scales would not be identical. Moreover, due to the unequal scale of $\widetilde{\bm{\theta}}$, it is not possible to derive the limited null distribution of $||\widetilde{\bm{\theta}}||_{\infty}$ even under weak correlation assumption. In this paper, we first provide the Bahadur representation and Gaussian approximation of the location estimator proposed in Feng et al. (2016). We then propose a new max-type test statistic and establish its limit null distribution under some mild conditions.
2.1 Bahadur representation and Gaussian approximation
Motivated by Feng et al. (2016), we suggest to find a pair of diagonal matrix $\mathbf{D}$ and vector $\bm{\theta}$ for each sample that simultaneously satisfy
$\frac{1}{n}\sum_{i=1}^{n}U\left(\bm{\epsilon}_{i}\right)=0\text{ and }\frac{p}{n}\operatorname{diag}\left\{\sum_{i=1}^{n}U\left(\bm{\epsilon}_{i}\right)U\left(\bm{\epsilon}_{i}\right)^{T}\right\}=\mathbf{I}_{p},$ (2.4)
where $\bm{\epsilon}_{i}=\mathbf{D}^{-1/2}\left(\bm{X}_{i}-\bm{\theta}\right)$. ($\mathbf{D}$, $\left.\bm{\theta}\right)$ can be viewed as a simplified version of HettmanspergerRandles (HR) estimator without considering the off-diagonal elements of $\mathbf{S}$. We can adapt the recursive algorithm of Feng et al. (2016) to solve Equation 2.4. That is, repeat the following three steps until convergence:
(i) $\bm{\epsilon}_{i}\leftarrow\mathbf{D}^{-1/2}\left(\mathbf{X}_{i}-\bm{\theta}\right),\quad j=1,\cdots,n_{i}$;
(ii) $\bm{\theta}\leftarrow\bm{\theta}+\frac{\mathbf{D}^{1/2}\sum_{j=1}^{n}U\left(\bm{\epsilon}_{i}\right)}{\sum_{j=1}^{n}\left\|\epsilon_{i}\right\|^{-1}}$;
(iii) $\mathbf{D}\leftarrow p\mathbf{D}^{1/2}\operatorname{diag}\left\{n^{-1}\sum_{j=1}^{n}U\left(\bm{\epsilon}_{i}\right)U\left(\bm{\epsilon}_{i}\right)^{T}\right\}\mathbf{D}^{1/2}$.
The resulting estimators of location and diagonal matrix are denoted as $\hat{\bm{\theta}}$ and $\hat{\mathbf{D}}$. The sample mean and sample variances can be utilized as initial estimators. Regrettably, no evidence has been found to confirm the convergence of the aforementioned algorithm, even in low-dimensional scenarios, despite its consistent practical effectiveness. The existence or uniqueness of the HR estimate mentioned above also lacks proof. This topic certainly warrants further investigation.
In this section, we investigate some theoretical properties based on maximum-norm about $\hat{\bm{\theta}}$. Similar with the proof of Lemma 1 and Theorem 1 in Cheng et al. (2023), we give the Bahadur representation of $\hat{\bf D}^{-1/2}(\hat{\bm{\theta}}-\bm{\theta})$ and the Gaussian approximation result for $\hat{\bf D}^{-1/2}(\hat{\bm{\theta}}-\bm{\theta})$ over hyperrectangles. Based on Gaussian approximation, we can easily derive the limiting distribution of $\hat{\bm{\theta}}$ based on the maximum-norm.
For $i=1,2,\cdots,n$, we denote $\bm{U}_{i}=U(\mathbf{D}^{-1/2}(\bm{X}_{i}-\bm{\theta}))$ and $R_{i}=\|\mathbf{D}^{-1/2}(\bm{X}_{i}-\bm{\theta})\|$ as the scale-invariant spatial-sign and radius of $\bm{X}_{i}-\bm{\theta}$, respectively. The moments of $R_{i}$ is defined as $\zeta_{k}=\mathbb{E}\left(R_{i}^{-k}\right)$ for $k=1,2,3,4$. Denote $\bm{W}_{i}=\left(W_{i,1},\ldots,W_{i,p}\right)^{\top}$, the assumption is as follows.
Assumption 1.
$W_{i,1},\ldots,W_{i,p}$ are i.i.d. symmetric random variables with $\mathbb{E}\left(W_{i,j}\right)=0,\mathbb{E}\left(W_{i,j}^{2}\right)=$ 1 , and $\left\|W_{i,j}\right\|_{\psi_{\alpha}}\leqslant c_{0}$ with some constant $c_{0}>0$ and $1\leqslant\alpha\leqslant 2$.
Assumption 2.
The moments $\zeta_{k}=\mathbb{E}\left(R_{i}^{-k}\right)$ for $k=1,2,3,4$ exist for large enough $p$. In addition, there exist two positive constants $\underline{b}$ and $\bar{B}$ such that $\underline{b}\leqslant\lim\sup_{p}\mathbb{E}\left(R_{i}/\sqrt{p}\right)^{-k}\leqslant\bar{B}$ for $k=1,2,3,4$.
Assumption 3.
The shape matrix ${\bf R}=\mathbf{D}^{-1/2}\Gamma\Gamma^{\top}\mathbf{D}^{-1/2}=\left(\sigma_{j\ell}\right)_{p\times p}$ satisfies $\max_{j=1,\cdots,p}\sum_{\ell=1}^{p}\left|\sigma_{j\ell}\right|\leqslant a_{0}(p)$. In addition, $\lim\inf_{p\rightarrow\infty}\min_{j=1,2,\cdots,p}d_{j}>\underline{d}$ for some constant $\underline{d}>0$, where $\mathbf{D}=\operatorname{diag}\{d_{1}^{2},d_{2}^{2},\cdots,d_{p}^{2}\}$.
Remark 1  Assumption 1 is the same as Condition C.1 in Cheng et al. (2023), which ensure that $\bm{\theta}$ in model (2.1) is the population spatial median and $W_{i,j}$ has a sub-exponential distribution. If $\bm{W}_{i}\sim N(\bm{0},{\bf I}_{p})$, $\bm{X}_{i}$ follows a elliptical symmetric distribution. Assumption 2 extend the Assumption 1 in Zou et al. (2014), which indicates that $\zeta_{k}\asymp p^{-k/2}$. Assumption 3 means the correlation between those variables could not be too large, which is similar to the matrix class in Bickel and Levina (2008).
The following lemma shows a Bahadur representation of $\hat{\bm{\theta}}$, which is the basis of Gaussian approximation result in Theorem 1.
Lemma 1.
(Bahadur representation) Under Assumption 1-3 with $a_{0}(p)\asymp p^{1-\delta}$, if $\log p=o(n^{1/3})$ and $\log n=o(p^{1/3\wedge\delta})$, we have that:
$n^{1/2}\hat{\mathbf{D}}^{-1/2}(\hat{\bm{\theta}}-\bm{\theta})=n^{-1/2}\zeta_{1}^{-1}\sum_{i=1}^{n}\bm{U}_{i}+C_{n},$
where
$\|C_{n}\|_{\infty}=O_{p}(n^{-1/4}\log^{1/2}(np)+p^{-(1/6\wedge\delta/2)}\log^{1/2}(np)+n^{-1/2}(\log p)^{1/2}\log^{1/2}(np)).$
Remark 2  Feng et al. (2016) derived the Bahadur representation of the estimator $\hat{\bm{\theta}}$, where the remainder term $||C_{n}||$ is $o_{p}(\zeta_{1}^{-1})$, assuming a symmetric elliptical distribution. In this context, we provide the rate of the remainder term $C_{n}$ subject to a maximum-norm constraint. It’s important to note that in this Lemma, we scale the spatial-median estimator $\hat{\bm{\theta}}$ by $\hat{\bf D}^{-1/2}$. This is a departure from much of the existing literature on the Bahadur representation of the spatial median, which does not exhibit scalar invariance. Such works include Zou et al. (2014), Cheng et al. (2019), Li and Xu (2022), and Cheng et al. (2023).
Let $\mathcal{A}^{\text{re }}=\left\{\prod_{j=1}^{p}\left[a_{j},b_{j}\right]:-\infty\leqslant a_{j}\leqslant b_{j}\leqslant\infty,j=1,\ldots,p\right\}$ be the class of rectangles in $\mathbb{R}^{p}$. Based on the Bahadur representation of $\hat{\bm{\theta}}_{n}$, we acquire the following Gaussian approximation of $\hat{\mathbf{D}}^{-1/2}\left(\hat{\bm{\theta}}-\bm{\theta}\right)$ in rectangle $\mathcal{A}^{\text{re }}$.
Lemma 2.
(Gaussian approximation) Assume Assumption 1-3 with $a_{0}(p)\asymp p^{1-\delta}$ for some positive constant $\delta\leqslant 1/2$ hold. If $\log p=o\left(n^{1/5}\right)$ and $\log n=o\left(p^{1/3\wedge\delta}\right)$, then
$\rho_{n}\left(\mathcal{A}^{\mathrm{re}}\right)=\sup_{A\in\mathcal{A}^{\mathrm{re}}}\left|\mathbb{P}\left\{n^{1/2}\hat{\mathbf{D}}^{-1/2}\left(\hat{\bm{\theta}}-\bm{\theta}\right)\in A\right\}-\mathbb{P}(\bm{G}\in A)\right|\rightarrow 0,$
as $n\rightarrow\infty$, where $\bm{G}\sim N\left(0,\zeta_{1}^{-2}\mathbf{\Sigma}_{u}\right)$ with $\mathbf{\Sigma}_{u}=\mathbb{E}\left(\bm{U}_{1}\bm{U}_{1}^{\top}\right)$.
The Gaussian approximation for $\hat{\bm{\theta}}$ indicates that the probabilities $\mathbb{P}\left\{n^{1/2}\hat{\mathbf{D}}^{-1/2}\left(\hat{\bm{\theta}}-\bm{\theta}\right)\in A\right\}$ can be approximated by that of a centered Gaussian random vector with covariance matrix $\zeta_{1}^{-2}\mathbf{\Sigma}_{u}$ for hyperrectangles $A\in\mathcal{A}^{\text{re }}$.
Since the region $\mathcal{A}^{t}=\left\{\prod_{j=1}^{p}\left[a_{j},b_{j}\right]:-\infty=a_{j}\leqslant b_{j}=t,j=1,\ldots,p\right\}$ used in the following corollary is contained in the set $\mathcal{A}^{\mathrm{re}}$, it is clear that the Corollary 1 follows.
Corollary 1.
Suppose the conditions assumed in Lemma 2 hold, as $n\rightarrow\infty$,
$\rho_{n}=\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(n^{1/2}\|\hat{\mathbf{D}}^{-1/2}(\hat{\bm{\theta}}-\bm{\theta})\|_{\infty}\leqslant t\right)-\mathbb{P}\left(\|\bm{G}\|_{\infty}\leqslant t\right)\right|\rightarrow 0.$
where $\bm{G}\sim N\left(0,\zeta_{1}^{-2}\mathbf{\Sigma}_{u}\right)$
Taking into account the relationships between $\mathbf{\Sigma}_{u}$ and $p^{-1}{\bf R}$, we propose a more straightforward Gaussian approximation.
Lemma 3.
(Variance approximation) Suppose $\bm{G}\sim N\left(0,\zeta_{1}^{-2}\mathbf{\Sigma}_{u}\right)$ and $\bm{Z}\sim N\left(0,\zeta_{1}^{-2}p^{-1}\mathbf{R}\right)$, as $(n,p)\rightarrow\infty$,
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(\|\bm{Z}\|_{\infty}\leqslant t\right)-\mathbb{P}\left(\|\bm{G}\|_{\infty}\leqslant t\right)\right|\rightarrow 0.$
By integrating Corollary 1 and Lemma 3, we can readily derive the principal theorem of Gaussian approximation.
Theorem 1.
Assume Assumptions 1-3 with $a_{0}(p)=p^{1-\delta}$ for some positive constant $\delta\leqslant 1/2$ hold. If $\log p=o\left(n^{1/5}\right)$ and $\log n=o\left(p^{1/3\wedge\delta}\right)$, then
$\widetilde{\rho}_{n}=\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(n^{1/2}\|\hat{\mathbf{D}}^{-1/2}(\hat{\bm{\theta}}-\bm{\theta})\|_{\infty}\leqslant t\right)-\mathbb{P}\left(\|\bm{Z}\|_{\infty}\leqslant t\right)\right|\rightarrow 0,$
where $\bm{Z}\sim N\left(0,\zeta_{1}^{-2}p^{-1}{\bf R}\right)$.
2.2 Max-type test procedure
In order to guarantee that the maximum value of a sequence of normal variables adheres to a Gumbel limiting distribution, we introduce Assumption 4. This assumption is employed to specify the necessary correlation among variables.
Assumption 4.
Let ${\bf R}=\left(\sigma_{ij}\right)_{1\leq i,j\leq p}$. For some $\varrho\in(0,1)$, assume $|\sigma_{ij}|\leq\varrho$ for all $1\leq i<j\leq p$ and $p\geq 2$. Suppose $\left\{\delta_{p};p\geq 1\right\}$ and $\left\{\kappa_{p};p\geq 1\right\}$ are positive constants with $\delta_{p}=o(1/\log p)$ and $\kappa=\kappa_{p}\rightarrow 0$ as $p\rightarrow\infty$. For $1\leq i\leq p$, define $B_{p,i}=\left\{1\leq j\leq p;\left|\sigma_{ij}\right|\geq\delta_{p}\right\}$ and $C_{p}=\left\{1\leq i\leq p;\left|B_{p,i}\right|\geq p^{\kappa}\right\}$. We assume that $\left|C_{p}\right|/p\rightarrow 0$ as $p\rightarrow\infty$.
Remark 3  Assumption 4 aligns with Assumption (2.2) in Feng et al. (2022a). This assumption stipulates that for each variable, the count of other variables that exhibit a strong correlation with it cannot be excessively large. To the best of our understanding, this is the least restrictive assumption in the literature that allows for the limiting null distribution of the maximum of correlated normal random variables to follow a Gumbel distribution. Both Assumption 3 and 4 pertain to the correlation matrix ${\bf R}$. We examine two specific cases that satisfy both of these conditions. The first case is the classic AR(1) structure, denoted as ${\bf R}=(\rho^{|i-j|})_{1\leq i,j\leq p},\rho\in(-1,1)$. In this scenario, $\sum_{l=1}^{p}|\sigma_{jl}|\to\frac{1}{1+\rho}$, which allows $\delta$ to be one in Assumption 3. For Assumption 4, we set $\delta_{p}=(\log p)^{-2}$, leading to $B_{p,i}=\{j:|i-j|\leq-2\log\log p/\log|\rho|\}$. As a result, $|B_{p,i}|\leq-4\log\log p/\log|\rho|<p^{\kappa}$ with $\kappa=5\log\log p/\log p$, which implies $|C_{p}|=0$ and Assumption 4 is satisfied. The second case involves a banded correlation matrix, where $\sigma_{ij}=0$ if $|i-j|>\ell$. Here, $\sum_{j=1}^{p}|\sigma_{ij}|=O(\ell)$ and $|B_{p,i}|\leq\ell$ for $\delta_{p}=(\log p)^{-2}$. Therefore, Assumptions 3 and 4 will hold if $\ell=o(p^{\kappa})$ for any $\kappa\to 0$.
Suppose Assumption 1-4 holds, by the Theorem 2 in Feng et al. (2022a), we can see that $p\zeta_{1}^{2}\max_{1\leq i\leq p}Z_{i}^{2}-2\log p+\log\log p$ converges to a Gumbel distribution with cdf $F(x)=\exp\left\{-\frac{1}{\sqrt{\pi}}e^{-x/2}\right\}$ as $p\rightarrow\infty$. In combining with the Theorem 1 we can conclude that,
$\mathbb{P}\left(n^{1/2}\left\|\hat{\mathbf{D}}^{-1/2}(\hat{\bm{\theta}}-\bm{\theta})\right\|_{\infty}^{2}p\zeta_{1}^{2}-2\log p+\log\log p\leq x\right)\rightarrow\exp\left\{-\frac{1}{\sqrt{\pi}}e^{-x/2}\right\}.$ (2.5)
Next we replace $E\left(R^{-1}\right)$ with its estimators. Denote $\hat{R}_{i}=\left\|\hat{\mathbf{D}}^{-1/2}\left(\mathbf{X}_{i}-\hat{\bm{\theta}}\right)\right\|$. Then the estimator is defined as $\hat{E\left(R_{i}^{-1}\right)}=\frac{1}{n}\sum_{i=1}^{n}\hat{R}_{i}^{-1}:=\hat{\zeta}_{1}$, and the proof of consistency is shown in Lemma 7. Because the convergence rate of maximum is very slow, we propose a adjust max-type test statistic which based on the scalar-invariant spatial median $\hat{\bm{\theta}}$,
$T_{MAX}=n\left\|\hat{\mathbf{D}}^{-1/2}\hat{\bm{\theta}}\right\|_{\infty}^{2}\hat{\zeta}_{1}^{2}p\cdot\left(1-n^{-1/2}\right).$
Theorem 2.
Suppose the conditions in Theorem 1 and Assumption 4 hold. Under the null hypothesis, as $(n,p)\to\infty$, we have
$P\left(T_{MAX}-2\log p+\log\log p\leq x\right)\rightarrow\exp\left\{-\frac{1}{\sqrt{\pi}}e^{-x/2}\right\}.$
According to Theorem 2, a level- $\alpha$ test will then be performed through rejecting $H_{0}$ when $T_{MAX}-2\log p+\log\log p$ is larger than the $(1-\alpha)$ quantile $q_{1-\alpha}=-\log\pi-2\log\log(1-\alpha)^{-1}$ of the Gumbel distribution $F(x)$.
The following theorem demonstrates the consistency of the proposed max-type test.
Theorem 3.
Suppose the conditions assumed in Theorem 2 holds. For any given $\alpha\in(0,1)$, if $\|\bm{\theta}\|_{\infty}\geq\widetilde{C}n^{-1/2}\{\log p-2\log\log(1-\alpha)^{-1}\}^{1/2}$, for some large enough constant $\widetilde{C}$, then
$\mathbb{P}(T_{MAX}-2\log p+\log\log p>q_{1-\alpha}\mid H_{1})\rightarrow 1,$
as $n\rightarrow\infty$.
Given a fixed significance level $\alpha$, the test $T_{MAX}$ attains consistency if $\|\bm{\theta}\|_{\infty}\geq\widetilde{C}\sqrt{\log p/n}$, provided that $\widetilde{C}$ is sufficiently large. This is the minimax rate optimal for testing against sparse alternatives, as stated in Theorem 3 of Cai et al. (2013). If $\widetilde{C}$ is adequately small, then it becomes impossible for any $\alpha$-level test to reject the null hypothesis with a probability approaching one. Therefore, Theorem 3 also demonstrates the optimality of our proposed test $T_{MAX}$.
To show the high dimensional asymptotic relative efficiency, we consider a special alternative hypothesis:
$H_{1}:\bm{\theta}=(\theta_{1},0,\cdots,0)^{T},\theta_{1}>0$
which means there are only one variable has nonzero means. Let $x_{\alpha}=2\log p-\log\log p+q_{1-\alpha}$. In this case,
$\mathbb{P}\left(\hat{d}_{1}^{-2}\hat{\theta}^{2}_{1}n\hat{\zeta}_{1}^{2}p\geq x_{\alpha}\right)\leq\mathbb{P}\left(T_{MAX}\geq x_{\alpha}\right)\leq\mathbb{P}\left(\hat{d}_{1}^{-2}\hat{\theta}^{2}_{1}n\hat{\zeta}_{1}^{2}p\geq x_{\alpha}\right)+\mathbb{P}\left(\max_{2\leq i\leq p}\hat{d}_{i}^{-2}\hat{\theta}^{2}_{i}n\hat{\zeta}_{1}^{2}p\geq x_{\alpha}\right).$
Under this special alternative hypothesis, we can easily have
$\mathbb{P}\left(\max_{2\leq i\leq p}\hat{d}_{i}^{-2}\hat{\theta}^{2}_{i}n\hat{\zeta}_{1}^{2}p\geq x_{\alpha}\right)\to\alpha,$
and
$\mathbb{P}\left(\hat{d}_{1}^{-2}\hat{\theta}^{2}_{1}n\hat{\zeta}_{1}^{2}p\geq x_{\alpha}\right)\to\Phi\left(-\sqrt{x_{\alpha}}+(np)^{1/2}d_{1}^{-1}\theta_{1}\zeta_{1}\right).$
So the power function of our proposed $T_{MAX}$ test is
$\beta_{MAX}(\bm{\theta})\in(\Phi\left(-\sqrt{x_{\alpha}}+(np)^{1/2}d_{1}^{-1}\theta_{1}\zeta_{1}\right),\Phi\left(-\sqrt{x_{\alpha}}+(np)^{1/2}d_{1}^{-1}\theta_{1}\zeta_{1}\right)+\alpha).$
Similarly, the power function of Cai et al. (2013)’s test is
$\beta_{CLX}(\bm{\theta})\in(\Phi\left(-\sqrt{x_{\alpha}}+n^{1/2}\varsigma_{1}^{-1}\theta_{1}\right),\Phi\left(-\sqrt{x_{\alpha}}+n^{1/2}\varsigma_{1}^{-1}\theta_{1}\right)+\alpha),$
where $\varsigma^{2}_{i}$ is the variance of $X_{ki},i=1,\cdots,p$. Thus, the asymptotic relative efficiency of $T_{MAX}$ with respective to Cai et al. (2013)’s test could be approximated as
$ARE(T_{MAX},T_{CLX})=\{E(R_{i}^{-1})\}^{2}E(R_{i}^{2}),$
which is the same as the asymptotic relative efficiency of Feng and Sun (2016)’s test with respective to Srivastava (2009)’s test. If $\bm{X}_{i}$ are generated from standard multivariate $t$-distribution with $\nu$ degrees of freedom $(\nu>2)$,
$ARE(T_{MAX},T_{CLX})=\frac{2}{\nu-2}\left(\frac{\Gamma((\nu+1)/2)}{\Gamma(\nu/2)}\right)^{2}.$
For different $\nu=3,4,5,6$, the above ARE are $2.54,1.76,1.51,1.38$, respectively. Under the multivariate normal distribution $(\nu=\infty)$, our $T_{MAX}$ test is the same powerful as Cai et al. (2013)’s test. However, our $T_{MAX}$ test is much more powerful than Cai et al. (2013)’s test under the heavy-tailed distributions.
3 Maxsum test
3.1 Exisiting Sum-type test
Feng et al. (2016) proposed the following sum-type test statistic:
$T_{SUM}=\frac{2}{n(n-1)}{\sum\sum}_{i<j}U\left(\hat{\mathbf{D}}_{ij}^{-1/2}\bm{X}_{i}\right)^{T}U\left(\hat{\mathbf{D}}_{ij}^{-1/2}\bm{X}_{j}\right)$
where $\hat{\mathbf{D}}_{ij}$ are the corresponding diagonal matrix estimator using leave-two-out sample $\left\{{\bm{X}}_{k}\right\}_{k\neq i,j}^{n}$.
By Feng et al. (2016), we have the following theorem and assumptions:
Assumption 5.
Variables $\left\{{\bm{X}}_{1},\ldots,{\bm{X}}_{n}\right\}$ in the $n$-th row are independently and identically distributed (i.i.d.) from $p$-variate elliptical distribution with density functions $\operatorname{det}(\bm{\Sigma})^{-1/2}\cdot$ $g\left(\left\|\bm{\Sigma}^{-1/2}(\mathbf{x}-\bm{\theta})\right\|\right)$ where $\bm{\theta}$ ’s are the symmetry centers and $\bm{\Sigma}$ ’s are the positive definite symmetric $p\times p$ scatter matrices.
Assumption 6.
$\operatorname{tr}\left({\bf R}^{4}\right)=o\left(\operatorname{tr}^{2}\left({\bf R}^{2}\right)\right)$.
Assumption 7.
(i) $\operatorname{tr}\left({\bf R}^{2}\right)-p=o\left(n^{-1}p^{2}\right)$, (ii) $n^{-2}p^{2}/\operatorname{tr}\left({\bf R}^{2}\right)=O(1)$ and $\log p=o(n)$.
The following Lemma restate the Theorem 1 in Feng and Sun (2016), which gives the asymptotic null distribution of $T_{SUM}$ under the symmetric elliptical distribution assumption.
Lemma 4.
Under Assumptions 5-7. and $H_{0}$, as $\min\left(p,n\right)\rightarrow\infty,T_{SUM}/\sigma_{n}\stackrel{{\scriptstyle d}}{{\rightarrow}}N(0,1)$, where $\sigma_{n}^{2}=\frac{2}{n(n-1)p^{2}}\operatorname{tr}\left({{\bf R}}^{2}\right)$.
To broaden the application, we re-derive the limiting null distribution of $T_{SUM}$ under a more generalized model (2.1).
Theorem 4.
Under Assumptions 1-3,6-7. and $H_{0}$, as $\min\left(p,n\right)\rightarrow\infty,T_{SUM}/\sigma_{n}\stackrel{{\scriptstyle d}}{{\rightarrow}}N(0,1)$.
Similar to Feng and Sun (2016), we propose the following estimator to estimate the trace term in $\sigma_{n}^{2}$
$\widehat{\operatorname{tr}\left(\mathbf{R}^{2}\right)}=\frac{p^{2}}{n(n-1)}\sum_{i=1}^{n}\sum_{j\neq i}^{n}\left(U\left(\hat{\mathbf{D}}_{ij}^{-1/2}\left(\bm{X}_{i}-\hat{\bm{\theta}}_{ij}\right)\right)^{T}U\left(\hat{\mathbf{D}}_{ij}^{-1/2}\left(\bm{X}_{j}-\hat{\bm{\theta}}_{ij}\right)\right)\right)^{2},$
[…]
And we also re-derive the asymptotic distribution of $T_{SUM}$ under the following alternative hypothesis:
$H_{1}:\bm{\theta}^{\top}\mathbf{D}^{-1}\bm{\theta}=O\left(\zeta_{1}^{-2}\sigma\right)\text{ and }\bm{\theta}^{\top}{\bf R}\bm{\theta}=o\left(\zeta_{1}^{-2}np\sigma^{2}\right)$ (3.1)
Theorem 5.
Under Assumptions 1-3,6-7 and the alternative hypothesis (3.1), as $\min\left(p,n\right)\rightarrow\infty$,
$\frac{T_{SUM}-\zeta_{1}^{2}\bm{\theta}^{\top}\mathbf{D}^{-1}\bm{\theta}}{\sigma_{n}}\stackrel{{\scriptstyle d}}{{\rightarrow}}N(0,1).$
By Theorem 4 and 5, the asymptotic power function of $T_{SUM}$ is
$\beta_{SUM}\left(\bm{\theta}\right)=\Phi\left(-z_{\alpha}+\frac{\zeta_{1}^{2}np\bm{\theta}^{T}\mathbf{D}^{-1}\bm{\theta}}{\sqrt{2\operatorname{tr}\left(\mathbf{R}^{2}\right)}}\right).$
Hence, $T_{SUM}$ is expected to perform well under the dense alternative hypothesis. For a more detailed discussion on the asymptotic relative efficiency of $T_{SUM}$ compared to other tests, refer to Feng and Sun (2016). The power comparison between $T_{SUM}$ and $T_{MAX}$ will be addressed in the following subsection.
3.2 Maxsum test
In this subsection, we demonstrate that our proposed test statistic $T_{MAX}$ is asymptotically independent of the statistic $T_{SUM}$ presented in Feng and Sun (2016). This allows us to carry out a Cauchy $p$-value combination of the two asymptotically independent $p$-values, resulting in a new test. This test is tailored to accommodate both sparse and dense alternatives.
Assumption 8.
There exist $C>0$ and $\varrho\in(0,1)$ so that $\max_{1\leq i<j\leq p}\left|\sigma_{ij}\right|\leq\varrho$ and $\max_{1\leq i\leq p}\sum_{j=1}^{p}\sigma_{ij}^{2}\leq(\log p)^{C}$ for all $p\geq 3;p^{-1/2}(\log p)^{C}\ll\lambda_{\min}({{\bf R}})\leq\lambda_{\max}({\bf R})\ll\sqrt{p}(\log p)^{-1}$ and $\lambda_{\max}({\bf R})/\lambda_{\min}({\bf R})=O\left(p^{\tau}\right)$ for some $\tau\in(0,1/4)$.
Remark 4  Assumption 8 is the same as the condition (2.3) in Feng et al. (2022b). As shown in Feng et al. (2022b), Assumption 8 is more restrictive than Assumption 3, 4 and 6. Under Assumption 8, we have $p^{1/2}(\log p)^{C}\lesssim{\rm tr}({\bf R}^{2})\lesssim p^{3/2}\log^{-1}p$. So Assumption 7 will hold if $n=o(p^{3/2}\log p)$ and $p^{3/4}(\log p)^{-C/2}=O(n)$. Intuitively speaking, if all the eigenvalues of ${\bf R}$ are bounded and $p/n\to c\in(0,\infty)$, all the assumptions 3, 4, 6, 7 and 8 hold.
Theorem 6.
Under Assumptions 1-4, 7-8 and $H_{0}$, $T_{SUM}$ and $T_{MAX}$ are asymptotically independent, i.e.
$\mathbb{P}\left(T_{SUM}/\sigma_{n}\leq x,T_{MAX}-2\log p+\log\log p\leq y\right)\to\Phi(x)F(y),$
as $n,p\to\infty$.
According to Theorem 6, we suggest combining the corresponding $p$-values by using Cauchy Combination Method (Liu and Xie, 2020), to wit,
$p_{CC}$ $=1-G[0.5\tan\{(0.5-p_{MAX})\pi\}+0.5\tan\{(0.5-p_{SUM})\pi\}],$
$p_{MAX}$ $=1-F(T_{MAX}-2\log p+\log\log p),$
$p_{SUM}$ $=1-\Phi(T_{SUM}/\hat{\sigma}_{n}),$
where $G(\cdot)$ is the CDF of the the standard Cauchy distribution. If the final $p$-value is less than some pre-specified significant level $\alpha\in(0,1)$, then we reject $H_{0}$.
Next, we consider the relationship between $T_{SUM}$ and $T_{MAX}$ under local alternative hypotheses:
$H_{1}:\|\bm{\theta}\|=O(\zeta_{1}^{-2}\sigma),\|{\bf R}^{1/2}\bm{\theta}\|=o\left(\zeta_{1}^{-2}np\sigma^{2}\right)\text{ and }|\mathcal{A}|=o\left(\frac{\lambda_{\min}({\bf R})[\text{tr}({\bf R}^{2})]^{1/2}}{(\log p)^{C}}\right),$ (3.2)
where $\mathcal{A}=\{i\mid\theta_{i}\not=0,1\leq i\leq p\}$, $\bm{\theta}=(\theta_{1},\theta_{2},\cdots,\theta_{p})^{\top}$. The following theorem establish the asymptotic independence between $T_{SUM}$ and $T_{MAX}$ under this special alternative hypothesis.
Theorem 7.
Under Assumptions 1-4, 6-8 and the alternative hypothesis (3.2), $T_{SUM}$ and $T_{MAX}$ are asymptotically independent, i.e.
$\mathbb{P}\left(T_{SUM}/\sigma_{n}\leq x,T_{MAX}-2\log p+\log\log p\leq y\right)\to$
$\mathbb{P}\left(T_{SUM}/\sigma_{n}\leq x\right)\mathbb{P}\left(T_{MAX}-2\log p+\log\log p\leq y\right),$
as $n,p\to\infty$.
According to Long et al. (2023), the Cauchy combination-based test has more power than the test based on the minimum of $p_{MAX}$ and $p_{SUM}$, which is also known as the minimal p-value combination. This is represented as $\beta_{M\wedge S,\alpha}=P(\min\{{\rm p}_{MAX},{\rm p}_{SUM}\}\leq 1-\sqrt{1-\alpha})$.
It is clear that:
$\beta_{M\wedge S,\alpha}$ $\geq P(\min\{{\rm p}_{MAX},{\rm p}_{SUM}\}\leq\alpha/2)$
$=\beta_{MAX,\alpha/2}+\beta_{SUM,\alpha/2}-P({\rm p}_{MAX}\leq\alpha/2,{\rm p}_{SUM}\leq\alpha/2)$
$\geq\max\{\beta_{MAX,\alpha/2},\beta_{SUM,\alpha/2}\}.$ (3.3)
On the other hand, under the local alternative hypothesis (3.2), we have:
$\beta_{M\wedge S,\alpha}\geq\beta_{MAX,\alpha/2}+\beta_{SUM,\alpha/2}-\beta_{MAX,\alpha/2}\beta_{SUM,\alpha/2}+o(1),$ (3.4)
which is due to the asymptotic independence implied by Theorem 7.
For a small $\alpha$, the difference between $\beta_{MAX,\alpha}$ and $\beta_{MAX,\alpha/2}$ is small, and the same applies to $\beta_{SUM,\alpha}$. Therefore, according to equations (3.2) and (3.4), the power of the adaptive test is at least as large as, or even significantly larger than, that of either the max-type or sum-type test. For a detailed comparison of the performance of each test type under varying conditions of sparsity and signal strength, please refer to Table 1 in Ma et al. (2024).
4 Simulation
In this section, we incorporated various methods into our study:
  • the proposed test $T_{MAX}$, referred as SS-MAX;
  • sum-type test proposed by Feng and Sun (2016), referred as SS-SUM;
  • the proposed test $T_{CC}$, referred as SS-CC;
  • max-type method proposed by Cai et al. (2013) , referred as MAX;
  • sum-type method proposed by Srivastava (2009), referred as SUM;
  • combination test proposed by Feng et al. (2022a), referred as COM.
The following scenarios are firstly considered.
  • (I)
    Multivariate normal distribution. ${\bf X}_{i}\sim N(\bm{\theta},{\bf\Sigma})$.
  • (II)
    Multivariate $t$-distribution $t_{p,4}$. ${\bf X}_{i}$’s are generated from $t_{p,4}$ with location parameter ${\bm{\theta}}$ and scatter matrix ${\bf\Sigma}$.
  • (III)
    Multivariate mixture normal distribution $\mbox{MN}_{p_{n},\gamma,9}$. ${\bf X}_{i}$’s are generated from $\gamma f_{p_{n}}(\bm{\theta},{\bf\Sigma})+(1-\gamma)f_{p_{n}}(\bm{\theta},9{\bf\Sigma})$, denoted by $\mbox{MN}_{p_{n},\gamma,9}$, where $f_{p_{n}}(\cdot;\cdot)$ is the density function of $p_{n}$-variate multivariate normal distribution. $\gamma$ is chosen to be 0.8.
Here we consider the scatter matrix ${\bf\Sigma}=(0.5^{|i-j|})_{1\leq i,j\leq p}$. Two sample sizes $n=50,100$ and three dimensions $p=200,400,600$ are considered. All the findings in this section are derived from 1000 repetitions. Table 1 presents the empirical sizes of the six tests mentioned above. It was observed that the spatial-sign based tests–SS-MAX, SS-SUM, and SS-CC–are able to effectively manage the empirical sizes in a majority of scenarios. However, when dealing with distributions that are not multivariate normal, the SUM test tends to have empirical sizes that fall below the nominal level. Similarly, the COM test also exhibits smaller sizes when operating under non-normal distributions.
To compare the power performance of each test, we consider $\bm{\theta}=(\kappa,\kappa,\kappa,0,\cdots,0)$ where the first $s$ components of $\bm{\theta}$ are all equal to $\kappa=\sqrt{0.5/s}$. Figure 1 illustrates the power curves for each test. In the case of the multivariate normal distribution, SS-SUM and SUM exhibit similar performance, aligning with the findings of Feng and Sun (2016). The spatial-sign based max-type test procedure, SS-MAX, is slightly less powerful than its mean-based counterpart, MAX. The two combination type test procedures demonstrate comparable performance in this scenario. However, when dealing with non-normal distributions, the spatial-sign based test procedures surpass the mean-based ones. Moreover, the newly proposed test, SS-CC, outperforms the others in most scenarios. In extremely sparse scenarios ($s<5$), SS-CC’s performance is akin to SS-MAX. In highly dense scenarios ($s>10$), SS-CC performs similarly to SS-SUM. However, when the signal is neither very sparse nor very dense, SS-CC proves to be the most effective among all test procedures. This underscores the superiority of our proposed max-sum procedures, not only in handling signal sparsity but also in dealing with heavy-tailed distributions.
Table 1: Empirical size comparison of various tests with a nominal level 5%.
$n=50$ $n=100$
$p$ 200 400 600 200 400 600
Multivariate Normal Distribution
SS-MAX 0.051 0.061 0.049 0.025 0.04 0.032
SS-SUM 0.061 0.056 0.041 0.06 0.059 0.068
SS-CC 0.071 0.065 0.048 0.057 0.056 0.056
MAX 0.095 0.125 0.116 0.052 0.081 0.072
SUM 0.076 0.086 0.054 0.069 0.064 0.081
COM 0.095 0.108 0.089 0.063 0.072 0.058
Multivariate $t_{3}$ Distribution
SS-MAX 0.063 0.062 0.063 0.061 0.063 0.058
SS-SUM 0.067 0.053 0.064 0.062 0.064 0.053
SS-CC 0.058 0.061 0.068 0.058 0.052 0.059
MAX 0.044 0.052 0.047 0.033 0.036 0.04
SUM 0.005 0.001 0.001 0.002 0.001 0
COM 0.021 0.03 0.027 0.019 0.014 0.019
Multivariate Mixture Normal Distribution
SS-MAX 0.056 0.061 0.07 0.037 0.037 0.044
SS-SUM 0.066 0.05 0.051 0.054 0.058 0.061
SS-CC 0.067 0.058 0.064 0.052 0.046 0.059
MAX 0.037 0.042 0.056 0.031 0.038 0.03
SUM 0.004 0 0 0.007 0.002 0
COM 0.021 0.019 0.028 0.013 0.02 0.01
Figure 1: Power of tests with different sparsity levels over $(n,p)=(100,200)$.
Next, we consider the power comparison of those tests under different signal strength. Here we consider three sparsity level $s=2,20,50$ and the signal parameter $\kappa=\sqrt{\delta/s}$. Figures 2 through 4 present the power curves for various testing methods under Scenarios I to III. As the signal strength increases, the power of all tests also increases. Despite the presence of heavy-tailed distributions, spatial-sign based testing methods continue to surpass those based on means. Among all the tests, the proposed CC test consistently delivers the best performance.
Figure 2: Power of tests with different signal strength for multivariate normal distribution over $(n,p)=(100,200)$.
Figure 3: Power of tests with different signal strength for multivariate $t_{3}$ distribution over $(n,p)=(100,200)$.
Figure 4: Power of tests with different signal strength for multivariate mixture normal distribution over $(n,p)=(100,200)$.
As shown in Feng and Sun (2016), for the sum-type test procedure, SS-SUM is more powerful than the non scalar-invariant test (Wang et al., 2015). Here we also compare our proposed test $T_{MAX}$ with Cheng et al. (2023)’s test (abbreviated as CPZ hereafter) to show the importance of scalar-invariant for max-type test procedure. We consider two scatter matrix case for $\bm{\Sigma}$: (i) $\bm{\Sigma}=(0.5^{|i-j|})_{1\leq i,j\leq p}$; (ii) $\bm{\Sigma}={\bf D}^{1/2}{\bf R}{\bf D}^{1/2},{\bf R}=(0.5^{|i-j|})_{1\leq i,j\leq p},{\bf D}=\hbox{diag}\{d_{1},\cdots,d_{p}\}$ where $d_{i}=1,i\leq p/2,d_{i}=3,i>p/2$. The other settings are all the same as above. Table 2 presents the empirical sizes of the SS-MAX and CPZ tests. Both tests are capable of controlling the empirical sizes in the majority of cases. Moreover, we conduct a power comparison of these two max-type tests under identical settings as previously mentioned, but with two distinct scatter matrix cases. Figures 5 and 6 depict the power curves of SS-MAX and CPZ under scatter matrix cases (i) and (ii), respectively. We observe that SS-MAX performs comparably to CPZ when all elements of the diagonal matrix of $\bm{\Sigma}$ are equal. However, SS-MAX exhibits greater power than CPZ when the elements of the diagonal matrix of the scatter matrix are unequal, underscoring the necessity of scalar-invariance.
Table 2: Empirical size comparison of SS-MAX and CPZ tests with a nominal level 5%.
$n=50$ $n=100$
$p$ 200 400 600 200 400 600
Scatter Matrix Case (i)
Multivariate Normal Distribution
SS-MAX 0.046 0.04 0.054 0.03 0.036 0.058
CPZ 0.076 0.054 0.076 0.056 0.08 0.088
Multivariate $t_{3}$ Distribution
SS-MAX 0.06 0.09 0.096 0.06 0.054 0.054
CPZ 0.06 0.064 0.06 0.052 0.07 0.038
Multivariate Mixture Normal Distribution
SS-MAX 0.072 0.06 0.072 0.038 0.052 0.038
CPZ 0.086 0.08 0.074 0.078 0.066 0.046
Scatter Matrix Case (ii)
Multivariate Normal Distribution
SS-MAX 0.038 0.054 0.032 0.044 0.044 0.032
CPZ 0.07 0.052 0.074 0.058 0.056 0.05
Multivariate $t_{3}$ Distribution
SS-MAX 0.064 0.08 0.068 0.058 0.068 0.064
CPZ 0.07 0.05 0.064 0.064 0.086 0.066
Multivariate Mixture Normal Distribution
SS-MAX 0.062 0.056 0.048 0.038 0.036 0.05
CPZ 0.068 0.05 0.05 0.084 0.052 0.052
Figure 5: Power of max-type tests with different signal strength for matrix case (i) over $(n,p)=(100,200)$.
Figure 6: Power of tests with different signal strength for matrix case (ii) over $(n,p)=(100,200)$.
5 Application
In this section, we utilize our methods to tackle a financial pricing problem. Our goal is to test whether the expected returns of all assets are equivalent to their respective risk-free returns. Let $X_{ij}\equiv R_{ij}-\textrm{rf}_{i}$ denote the excess return of the $j$th asset at time $i$ for $i=1,\cdots,n$ and $j=1,\cdots,p$, where $R_{ij}$ is the return on asset $j$ during period $i$ and $\textrm{rf}_{i}$ is the risk-free return rate of all asset during period $i$. We study the following pricing model
$X_{ij}=\mu_{j}+\xi_{ij},$ (5.1)
for $i=1,\cdots,n$ and $j=1,\cdots,p$, or, in vector form, ${\bf X}_{i}=\bm{\mu}+\bm{\xi}_{i}$, where ${\bf X}_{i}=(X_{i1},\dots,X_{ip})^{\top}$, $\bm{\mu}=(\mu_{1},\dots,\mu_{p})^{\top}$, and $\bm{\xi}_{i}=(\xi_{i1},\dots,\xi_{ip})^{\top}$ is the zero-mean error vector. We consider the following hypothesis
$H_{0}:\bm{\mu}=\bm{0}~{}~{}\text{versus}~{}~{}H_{1}:\bm{\mu}\not=\bm{0}.$
We examined the weekly return rates of stocks that are part of the S&P 500 index from January 2005 to November 2018. The weekly data were derived from the stock prices every Friday. Over time, the composition of the index changed and some stocks were introduced during this period. Therefore, we only considered a total of 424 stocks that were consistently included in the S&P 500 index throughout this period. We compiled a total of 716 weekly return rates for each stock during this period, excluding Fridays that were holidays. Given the possibility of autocorrelation in the weekly stock returns, we applied the Ljung-Box test Ljung and Box (1978) at a 0.05 level for zero autocorrelations to each stock. We retained 280 stocks for which the Ljung-Box test at a 0.05 level was not rejected. It’s important to note that if we had used all 424 stocks, there might be autocorrelation between observations, which would violate our assumption and necessitate further studies.
Figure 7 show the histogram of standard deviation of those 280 securities. We found that the variances of those assets are obviously not equal. So the scalar-invariant test procedure is preferred. Thus, We apply the above six test procedures–SS-SUM,SS-MAX,SS-CC,MAX,SUM,COM on the total samples. All the tests reject the null hypothesis significantly. To evaluate the performance of our proposed tests and other competing tests for both small and large sample sizes, we randomly sampled $n=52K$ observations from the 716 weekly returns, where $K$ ranges from 3 to 8. This experiment was repeated 1000 times for each $n$ value.
Table 3 presents the rejection rates of six tests. We discovered that spatial-sign based test procedures outperform mean-based test procedures. This is primarily due to the heavy-tailed nature of asset returns. Figure 8 displays Q-Q plots of the weekly return rates of some stocks in the S&P 500 index. We observed that all data deviate from a normal distribution and exhibit heavy tails. Additionally, sum-type tests perform better than max-type test procedures, mainly because the alternative is dense. Figure 9 illustrates the $t$-test statistic for each stock. We noticed that many $t$-test statistics are larger than 2, and most of them are positive. Among these tests, the SS-CC test performs the best. Although the SS-SUM outperforms the SS-MAX, the SS-MAX still retains some power in all cases. As indicated in the theoretical results in Subsection 3.2, our proposed Cauchy Combination would be more powerful than both max-type and sum-type tests in this scenario. Therefore, the application of real data also demonstrates the superiority of our proposed maxsum test procedure.
It’s worth noting that the rejection of the null hypothesis, which suggests that return rates are not solely composed of risk-free rates on average, aligns with the perspectives of numerous economists. Indeed, the consideration of a non-zero excess return rate and the attempt to model it has spurred a vast amount of research on factor pricing models in finance (Sharpe, 1964; Fama and French, 1993, 2015). These models, which have many practical applications, operate under the Arbitrage Pricing Theory (Ross, 1976). Recently, numerous studies have focused on the high-dimensional alpha test under the linear factor pricing model, including works by Fan et al. (2015); Pesaran and Yamagata (2017); Feng et al. (2022c); Liu et al. (2023). Notably, Liu et al. (2023) proposed a spatial-sign based sum-type test procedure for testing alpha for heavy-tailed distributions. It would be intriguing to extend the methods presented in this paper to propose a spatial sign based max-type and maxsum-type test procedures for testing alpha. This is an area that warrants further exploration.
Figure 7: Histogram of standard deviation of US securities.
Figure 8: Q-Q plots of the weekly return rates of some stocks with heavy-tailed distributions in the S&P500 index.
Table 3: The rejection rates of testing excess returns of the S&P stocks for $p=280$ and $n=52K$ with $K=3,\cdots,8$. For each $n$, we sampled 1000 data sets.
SS-MAX SS-SUM SS-CC MAX SUM COM
$n=156$ 0.295 0.361 0.380 0.124 0.219 0.204
$n=208$ 0.364 0.448 0.458 0.128 0.217 0.206
$n=260$ 0.424 0.542 0.556 0.140 0.276 0.246
$n=312$ 0.506 0.633 0.645 0.137 0.272 0.236
$n=364$ 0.652 0.738 0.758 0.143 0.317 0.282
$n=416$ 0.753 0.821 0.843 0.163 0.339 0.303
Figure 9: $t$ test statistics of the weekly excess return rates of each stock.
6 Conclusion
In this paper, we address a one-sample testing problem in high-dimensional settings for heavy-tailed distributions. We begin by providing a Bahadur representation and Gaussian approximation of the spatial median estimator, as discussed in Feng et al. (2016). Following this, we introduce a spatial-sign based max-type test procedure for sparse alternatives and establish the limit null distribution and consistency of the proposed max-type test statistic. Next, we reformulate the sum-type test statistic, originally proposed by Feng and Sun (2016), under a general model. This sum-type test exhibits superior performance under dense alternatives. Finally, we demonstrate the asymptotic independence between the aforementioned max-type test statistic and the sum-type test statistic, given some mild conditions. We then propose a Cauchy combination test procedure, which performs exceptionally well under both sparse and dense alternatives. Both simulation studies and real data applications underscore the superiority of the proposed maxsum-type test procedure.
We propose several directions for future research. Firstly, the sum-type test statistic in Feng and Sun (2016) only takes into account the direction of the sample, neglecting the information of the sample’s radius. Feng et al. (2021) introduced a more powerful inverse norm sign test. It would be intriguing to derive a max-type test statistic that also considers the radius of the sample. Furthermore, it remains an open question whether this new max-type test statistic maintains asymptotic independence with the sum-type test statistic proposed by Feng et al. (2021).
Secondly, the newly proposed methods can be extended to address other high-dimensional testing problems. These include the high-dimensional two-sample location problem (Chen and Qin, 2010; Feng et al., 2016), high-dimensional covariance matrix tests (Chen et al., 2010; Li and Chen, 2012; Cutting et al., 2017; Cheng et al., 2019), testing the martingale difference hypothesis in high dimension (Chang et al., 2023) and high-dimensional white noise test (Paindaveine and Verdebout, 2016; Chang et al., 2017; Feng et al., 2022d; Zhao et al., 2023). Additionally, the alpha test in the high-dimensional linear factor pricing model is a significant problem that has been explored in practical applications.
Thirdly, our paper’s theoretical results are predicated on the assumption of identical and independent distribution. However, there may occasionally be auto-correlations among the sample sizes. Recent literature, such as Zhang and Cheng (2018) and Chang et al. (2024), has considered the Gaussian approximation of the sample mean under a dependent assumption. Therefore, it would be intriguing to establish the Bahadur representation and Gaussian approximation of the spatial median in the context of dependent observations. Building on these findings, we can also suggest the implementation of max-type and maxsum-type testing methods for addressing high-dimensional location problem in the context of dependent observations (Ayyala et al., 2017; Ma et al., 2024).
7 Appendix
Recall that $\mathbf{D}=\operatorname{diag}\{d_{1}^{2},d_{2}^{2},\cdots,d_{p}^{2}\}$. For $i=1,2,\cdots,n$, $\bm{U}_{i}=U(\mathbf{D}^{-1/2}(\bm{X}_{i}-\bm{\theta}))$ and $R_{i}=\|\mathbf{D}^{-1/2}(\bm{X}_{i}-\bm{\theta})\|$ as the scale-invariant spatial-sign and radius of $\bm{X}_{i}-\bm{\theta}$, where $U(\bm{X})=\bm{X}/\|\bm{X}\|\mathbb{I}(\bm{X}\not=0)$ is the multivariate sign function of $\bm{X}$, with $\mathbb{I}(\cdot)$ being the indicator function. The moments of $R_{i}$ is defined as $\zeta_{k}=\mathbb{E}\left(R_{i}^{-k}\right)$ for k=1,2,3,4.
We denote the $\mathbf{D}$-estimated version $\bm{U}_{i}$ and $R_{i}$ as $\hat{R}_{i}=\|\hat{\bf D}^{-1/2}(\bm{X}_{i}-\bm{\theta})\|$ and $\hat{U}_{i}=\|\hat{\bf D}^{-1/2}(\bm{X}_{i}-\bm{\theta})/\|\hat{\bf D}^{-1/2}(\bm{X}_{i}-\bm{\theta})\|$, respectively, $i=1,2,\cdots,n$.
7.1 Proof of main lemmas
Lemma 5.
Under Assumption 1, we have $\mathbb{E}(U(\bm{W}_{i})^{\top}\mathbf{M}U(\bm{W}_{i}))^{2}=O(p^{-2}{\rm tr}(\mathbf{M}^{\top}\mathbf{M}))$.
Proof.  By Cauchy inequality and Assumption 1, we have
$\mathbb{E}U(\bm{W}_{i})_{l}^{2}U(\bm{W}_{i})_{k}^{2}\leq\frac{1}{p^{2}}\mathbb{E}\sum_{s=1}^{p}\sum_{t=1}^{p}U(\bm{W}_{i})_{s}^{2}U(\bm{W}_{i})_{t}^{2}=p^{-2}$
$\mathbb{E}U(\bm{W}_{i})_{l}^{4}\leq\frac{1}{p}\mathbb{E}\sum_{s=1}^{p}U(\bm{W}_{i})_{s}^{4}\leq\frac{1}{p}\mathbb{E}\sum_{s=1}^{p}\sum_{t=1}^{p}U(\bm{W}_{i})_{s}^{2}U(\bm{W}_{i})_{t}^{2}=p^{-1},$
and
$\mathbb{E}\left(U(\bm{W}_{i})_{l}U(\bm{W}_{i})_{k}U(\bm{W}_{i})_{s}U(\bm{W}_{i})_{t}\right)\leq\sqrt{E\left(U(\bm{W}_{i})_{l}^{2}U(\bm{W}_{i})_{k}^{2}\right)E\left(U(\bm{W}_{i})_{s}^{2}U(\bm{W}_{i})_{t}^{2}\right)}.$
By the Cauchy inequality,
$\sum_{i,k,s,t}a_{lk}a_{st}\leq\sqrt{\sum_{l,k}a_{lk}^{2}\sum_{s,t}a_{st}^{2}}\\ \leq\sqrt{\sum_{l,k}^{p}a_{lk}^{2}\sum_{s,t}^{p}a_{st}^{2}}=\text{tr}(\mathbf{M}^{\top}\mathbf{M}).$
Thus, we get
$E\left(\left(U(\bm{W}_{i})^{\top}\bf{M}U(\bm{W}_{i})\right)^{2}\right)$
$=$ $\sum_{l\not=k=1}^{p}\sum_{s\not=t=1}^{p}a_{lk}a_{st}\mathbb{E}\left(U(\bm{W}_{i})_{l}U(\bm{W}_{i})_{k}U(\bm{W}_{i})_{s}U(\bm{W}_{i})_{t}\right)+\sum_{l=1}^{p}\sum_{s=1}^{p}a_{ll}a_{ss}\mathbb{E}\left(U(\bm{W}_{i})_{l}^{2}U(\bm{W}_{i})_{s}^{2}\right)$
$\leq$ $p^{-2}\frac{p^{4}-p^{2}}{p^{4}}\text{tr}(\mathbf{M}^{\top}\mathbf{M})+p^{-1}\frac{p^{2}}{p^{4}}\text{tr}(\mathbf{M}^{\top}\mathbf{M})=O(p^{-2}\text{tr}(\bf M^{\top}\bf M)).$
Lemma 6.
Under Assumptions 1and 7, we have $\max_{j=1,2,\cdots p}(\hat{d}_{j}-d_{j})=O_{p}(n^{-1/2}(\log p)^{1/2})$.
Proof.  The proof of this lemma is along the same lines as the proof of Lemma A.2. in Feng et al. (2016), but differs in that the assumptions about the model in this paper are more general, with different constraints controlling the correlation matrix $R$.
Denote $\bm{\eta}=(\bm{\theta}^{\top},d_{1},d_{2},\cdots,d_{p})^{\top}$ and $\hat{\bm{\eta}}$ as the estimated version. By first-order Taylor expansion, we have
$U(\mathbf{D}^{-1/2}(\bm{X}_{i}-\bm{\theta}))$ $=\frac{\mathbf{D}^{-1/2}\mathbf{\Gamma}U(\bm{W}_{i})}{(1+U(\bm{W}_{i})^{\top}(\mathbf{R}-\mathbf{I}_{p})U(\bm{W}_{i}))^{1/2}}$ (7.1)
$=\mathbf{D}^{-1/2}\mathbf{\Gamma}U(\bm{W}_{i})+C_{1}U(\bm{W}_{i})^{\top}(\mathbf{R}-\mathbf{I}_{p})U(\bm{W}_{i})\mathbf{D}^{-1/2}\mathbf{\Gamma}U(\bm{W}_{i}),$
where $C_{1}$ is a bounded random variable between $-0.5$ and $-0.5(1+U(\bm{W}_{i})^{\top}(\mathbf{R}-\mathbf{I}_{p})U(\bm{W}_{i}))^{-3/2}$. By Cauchy inequality and Lemma 5 and Assumption 7, we get
$\mathbb{E}\left(U(\mathbf{D}^{-1/2}(\bm{X}_{i}-\bm{\theta})\right)$ $\leq C_{2}\left\{\mathbb{E}(U(\bm{W}_{i})^{\top}(\mathbf{R}-\mathbf{I}_{p})U(\bm{W}_{i}))^{2}\mathbb{E}(\mathbf{D}^{-1/2}(\bm{X}_{i}-\bm{\theta})^{2}\right\}^{1/2}$
$=O(p^{-1}\sqrt{\text{tr}(\mathbf{R}^{2})-p})=o(n^{-1/2}).$
Similarly, we can show that
$\mathbb{E}\left(\operatorname{diag}\left\{U(\mathbf{D}^{-1/2}(\bm{X}_{i}-\bm{\theta}))U(\mathbf{D}^{-1/2}(\bm{X}_{i}-\bm{\theta}))^{\top}\right\}-\frac{1}{p}\mathbf{I}_{p}\right)\leq O(n^{-1/2}),$
by first-order Taylor expansion for $U(\mathbf{D}^{-1/2}(\bm{X}_{i}-\bm{\theta}))U(\mathbf{D}^{-1/2}(\bm{X}_{i}-\bm{\theta}))^{\top}$, Cauchy inequality and Lemma 5. The above two equations define the functional equation for each component of $\bm{\eta}$,
$T_{j}(F,\eta_{j})=o_{p}(n^{-1/2}),$ (7.2)
where $F$ represent the distribution of $\bm{X}$, $\bm{\eta}=(\eta_{1},\cdots,\eta_{2p})$. Similar with Hettmansperger and Randles (2002), the linearisation of this equation shows
$n^{1/2}\left(\hat{\eta}_{j}-\eta_{j}\right)=-\mathbf{H}_{j}^{-1}n^{1/2}\left(T_{j}(F_{n},\eta_{j})-T_{j}(F,\eta_{j})\right)+o_{p}(1),$
where $F_{n}$ represents the empirical distribution function based on $\bm{X}_{1},\bm{X}_{2},\cdots,\bm{X}_{n}$, $\mathbf{H}_{j}$ is the corresponding Hessian matrix of the functional defined in Equation 7.2 and
$T(F_{n},\bm{\eta})=\left(n^{-1}\sum_{j=1}^{n}U(\mathbf{D}^{-1/2}(\bm{X}_{i}-\bm{\theta}))^{\top},\mathrm{vec}(\operatorname{diag}(n^{-1}U(\mathbf{D}^{-1/2}(\bm{X}_{i}-\bm{\theta}))U(\mathbf{D}^{-1/2}(\bm{X}_{i}-\bm{\theta}))^{\top}-\frac{1}{p}\mathbf{I}_{p}))\right).$
Thus, for each $\hat{d}_{j}$ we have
$\sqrt{n}(\hat{d}_{j}-d_{j})\stackrel{{\scriptstyle d}}{{\rightarrow}}N(0,\sigma^{2}_{d,j}).$
where $\sigma^{2}_{d,j}$ is the corresponding asymptotic variance. Define $\sigma_{d,max}=\max_{1\leq j\leq p}\sigma_{d,j}$. As $p\rightarrow\infty$,
$\mathbb{P}$ $(\max_{j=1,2,\cdots,p}(\hat{d}_{j}-d_{j})>\sqrt{2}\sigma_{d,max}n^{-1/2}(\log p)^{1/2})$
$\leq\sum_{j=1}^{p}\mathbb{P}(\sqrt{n}(\hat{d}_{j}-d_{j})>\sqrt{2}\sigma_{d,max}(\log p)^{1/2})$
$=\sum_{j=1}^{p}(1-\Phi(\sqrt{2}\sigma_{d,max}\sigma_{d,j}^{-1}(\log p)^{1/2}))\leq p(1-\Phi((2\log p)^{1/2}))$
$\leq\frac{p}{\sqrt{2\pi\log p}}\exp(-\log p)=(4\pi)^{-1/2}(\log p)^{-1/2}\rightarrow 0,$
which means that $\max_{j=1,2,\cdots p}(\hat{d}_{j}-d_{j})=O_{p}(n^{-1/2}(\log p)^{1/2})$.
Lemma 7.
Suppose the Assumptions in Lemma 6 hold, then $\hat{\zeta}_{1}\stackrel{{\scriptstyle p}}{{\rightarrow}}\zeta_{1}$.
Proof.  Denote $\hat{\bm{\mu}}=\hat{\bm{\theta}}-\bm{\theta}$.
$\|\hat{\mathbf{D}}^{-1/2}(\bm{X}_{i}-\hat{\bm{\theta}})\|=\|{\mathbf{D}}^{-1/2}(\bm{X}_{i}-{\bm{\theta}})\|($ $1+R_{i}^{-2}\|(\hat{\mathbf{D}}^{-1/2}-{\mathbf{D}}^{-1/2})(\bm{X}_{i}-{\bm{\theta}})\|^{2}$
$+R_{i}^{-2}\|\hat{\mathbf{D}}^{-1/2}\hat{\bm{\mu}}\|^{2}+2R_{i}^{-2}\bm{U}_{i}^{\top}(\hat{\mathbf{D}}^{-1/2}-{\mathbf{D}}^{-1/2}){\mathbf{D}}^{1/2}\bm{U}_{i})$
$-2R_{i}^{-1}\bm{U}_{i}^{\top}\hat{\mathbf{D}}^{-1/2}\hat{\bm{\mu}}-2R_{i}^{-1}\bm{U}_{i}\mathbf{D}^{1/2}(\hat{\mathbf{D}}^{-1/2}-{\mathbf{D}}^{-1/2})\hat{\mathbf{D}}^{-1/2}\hat{\bm{\mu}})^{1/2}.$
According to the proof and conclusion in Lemma 6, we can show that $R_{i}^{-2}\|(\hat{\mathbf{D}}^{-1/2}-\mathbf{D}^{-1/2})(\bm{X}_{i}-\bm{\theta})\|^{2}=O_{p}\left((\log p/n)^{1/2}\right)=o_{p}(1)$ and $R_{i}^{-2}\|\hat{\mathbf{D}}^{-1/2}\hat{\bm{\mu}}\|^{2}=O_{p}(n^{-1})=o_{p}(1)$ and by the Cauchy inequality, the other parts are also $o_{p}(1)$. So,
$n^{-1}\sum_{i=1}^{n}\left\|\hat{\mathbf{D}}^{-1/2}\left(\bm{X}_{i}-\hat{\bm{\theta}}\right)\right\|^{-1}=\left(n^{-1}\sum_{i=1}^{n}\left\|\mathbf{D}^{-1/2}\left(\bm{X}_{i}-\bm{\theta}\right)\right\|^{-1}\right)\left(1+o_{p}(1)\right).$
Obviously, $E\left(n^{-1}\sum_{i=1}^{n}R_{i}^{-1}\right)=\zeta_{i}$ and $\operatorname{var}\left(n^{-1}\zeta_{i}^{-1}\sum_{i=1}^{n}R_{i}^{-1}\right)=O\left(n^{-1}\right)$. Finally, the proof is completed.
Lemma 8.
Suppose Assumptions 1-3 holds with $a_{0}(p)\asymp p^{1-\delta}$ for some positive constant $\delta\leq 1/2$. Define a random $p\times p$ matrix $\hat{\mathbf{Q}}=n^{-1}\sum_{i=1}^{n}\hat{R}^{-1}_{i}\hat{\bm{U}}_{i}\hat{\bm{U}}_{i}^{\top}$ and let $\hat{\mathbf{Q}}_{jl}$ be the $(j,l)$th element of $\hat{\mathbf{Q}}$. Then,
$\left|\hat{\mathbf{Q}}_{j\ell}\right|\lesssim\zeta_{1}p^{-1}\left|\sigma_{j\ell}\right|+O_{p}\left(\zeta_{1}n^{-1/2}p^{-1}+\zeta_{1}p^{-7/6}+\zeta_{1}p^{-1-\delta/2}+\zeta_{1}n^{-1/2}(\log p)^{1/2}(p^{-5/2}+p^{-1-\delta/2})\right).$
Proof.  Recall that $\hat{\mathbf{Q}}=\frac{1}{n}\sum_{i=1}^{n}\hat{R}_{i}^{-1}\hat{\bm{U}}_{i}\hat{\bm{U}}_{i}^{\top}$, then,
$\hat{\mathbf{Q}}_{jl}=$ $\frac{1}{n}\sum_{i=1}^{n}\hat{R}_{i}^{-1}\hat{\bm{U}}_{i,j}\hat{\bm{U}}_{i,l}$
$=$ $\frac{1}{n}\sum_{i=1}^{n}v_{i}^{-1}(\hat{\mathbf{D}}^{-1/2}\mathbf{\Gamma}_{j}\bm{W}_{i})(\hat{\mathbf{D}}^{-1/2}\mathbf{\Gamma}_{l}\bm{W}_{i})^{\top}\|\hat{\mathbf{D}}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{-3}$
$=$ $\hat{\mathbf{D}}^{-1/2}\mathbf{D}^{1/2}\left\{\frac{1}{n}\sum_{i=1}^{n}v_{i}^{-1}(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{j}\bm{W}_{i})(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{l}\bm{W}_{i})^{\top}\|\hat{\mathbf{D}}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{-3}\right\}\hat{\mathbf{D}}^{-1/2}\mathbf{D}^{1/2}.$
We first consider the middle term,
$\frac{1}{n}\sum_{i=1}^{n}v_{i}^{-1}(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{j}\bm{W}_{i})(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{l}\bm{W}_{i})^{\top}\|\hat{\mathbf{D}}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{-3}$ (7.3)
$=$ $\frac{1}{n}\sum_{i=1}^{n}v_{i}^{-1}(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{j}\bm{W}_{i})(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{l}\bm{W}_{i})^{\top}\left\{\|\hat{\mathbf{D}}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{-3}-\|\mathbf{D}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{-3}\right\}$
$\quad+\frac{1}{n}\sum_{i=1}^{n}v_{i}^{-1}(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{j}\bm{W}_{i})(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{l}\bm{W}_{i})^{\top}\left\{\|\mathbf{D}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{-3}-p^{-3/2}\right\}$
$\quad\quad+n^{-1}p^{-3/2}\sum_{i=1}^{n}v_{i}^{-1}(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{j}\bm{W}_{i})(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{l}\bm{W}_{i})^{\top}.$
The first part in Equation 7.3:
$\mathbb{E}[\frac{1}{n}\sum_{i=1}^{n}v_{i}^{-1}(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{j}\bm{W}_{i})(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{l}\bm{W}_{i})^{\top}\left\{\|\hat{\mathbf{D}}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{-3}-\|\mathbf{D}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{-3}\right\}]$ (7.4)
$=$ $\mathbb{E}[v_{i}^{-1}(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{j}\bm{W}_{i})(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{l}\bm{W}_{i})^{\top}\left\{\|\hat{\mathbf{D}}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{-3}-\|\mathbf{D}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{-3}\right\}]$
$=$ $\mathbb{E}[v_{i}^{-1}(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{j}\bm{W}_{i})(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{l}\bm{W}_{i})^{\top}\left\{\|\hat{\mathbf{D}}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{-3}\|\mathbf{D}^{-1/2}\Gamma W_{i}\|^{-3}\left(\|\mathbf{D}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{3}-\|\hat{\mathbf{D}}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{3}\right)\right\}].$
To compute the order of Equation 7.4, we consider $\|\hat{\mathbf{D}}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{k}-\|\mathbf{D}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{k}$ for $k=1,2,\cdots$.
Firstly, for $k=2$, By the Lemma 6 and Assumption 3, we can see that,
$\max_{i=1,2,\cdots,p}({\frac{d_{i}}{\hat{d}_{i}}}-1)=\max_{i=1,2,\cdots,p}\frac{d_{i}-\hat{d}_{i}}{{\hat{d}_{i}}}=O_{p}(n^{-1/2}(\log p)^{1/2}).$
So, for $\|\hat{\mathbf{D}}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{2}$,
$\|\hat{\mathbf{D}}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{2}$
$=$ $\|(\hat{\mathbf{D}}^{-1/2}\mathbf{D}^{1/2}-\mathbf{I}_{p})\mathbf{D}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}+\mathbf{D}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{2}$
$=$ $\|\mathbf{D}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{2}+\|(\hat{\mathbf{D}}^{-1/2}\mathbf{D}^{1/2}-\mathbf{I}_{p})\mathbf{D}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{2}+\bm{W}_{i}\mathbf{\Gamma}^{\top}\mathbf{D}^{-1/2}(\hat{\mathbf{D}}^{-1/2}\mathbf{D}^{1/2}\mathbf{I}_{p})\mathbf{D}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}$
$\leq$ $\|\mathbf{D}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{2}\left(1+\max_{i=1,2,\cdots,p}({\frac{d_{i}}{\hat{d}_{i}}}-1)^{2}+\max_{i=1,2,\cdots,p}({\frac{d_{i}}{\hat{d}_{i}}}-1)\right)$
::=:= $\|\mathbf{D}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{2}\left(1+H\right),$
where $H:=\max_{i}({\frac{d_{i}}{\hat{d}_{i}}}-1)^{2}+\max_{i}({\frac{d_{i}}{\hat{d}_{i}}}-1)=O_{p}(n^{-1/2}(\log p)^{1/2})$.
For all integer $k$,
$\|\hat{\mathbf{D}}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{k}=$ $\|(\hat{\mathbf{D}}^{-1/2}\mathbf{D}^{1/2}-\mathbf{I}_{p})\mathbf{D}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}+\mathbf{D}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{k}$ (7.5)
$=$ $\left\{\|(\hat{\mathbf{D}}^{-1/2}\mathbf{D}^{1/2}-\mathbf{I}_{p})\mathbf{D}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}+\mathbf{D}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{2}\right\}^{k/2}$
$\leq$ $\|\mathbf{D}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{k}\left(1+H\right)^{k/2}$
::=:= $\|\mathbf{D}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{k}\left(1+H_{k}\right),$
where $H_{k}$ is defined as $H_{k}=(1+H)^{k/2}-1=O_{p}(n^{-1/2}(\log p)^{1/2})$.
Thus, from the proof of Lemma A3. in Cheng et al. (2023), Equation 7.4 equals
$\mathbb{E}[\frac{1}{n}\sum_{i=1}^{n}v_{i}^{-1}(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{j}\bm{W}_{i})(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{l}\bm{W}_{i})^{\top}\left\{\|\hat{\mathbf{D}}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{-3}-\|\mathbf{D}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{-3}\right\}]$
$=$ $\mathbb{E}[v_{i}^{-1}(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{j}\bm{W}_{i})(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{l}\bm{W}_{i})^{\top}\left\{\|\hat{\mathbf{D}}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{-3}-\|\mathbf{D}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{-3}\right\}]$
$=$ $\mathbb{E}[v_{i}^{-1}(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{j}\bm{W}_{i})(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{l}\bm{W}_{i})^{\top}\|\mathbf{D}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{-3}H_{3}]$
$=$ $\mathbb{E}[v_{i}^{-1}(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{j}\bm{W}_{i})(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{l}\bm{W}_{i})^{\top}(\|\mathbf{D}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{-3}-p^{-3/2})H_{3}]$
$\quad+p^{-3/2}\mathbb{E}[v_{i}^{-1}(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{j}\bm{W}_{i})(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{l}\bm{W}_{i})^{\top}H_{3}]$
$\lesssim$ $n^{-1/2}(\log p)^{1/2}\cdot\zeta_{1}p^{-5/2}(1+p^{3/2-\delta/2})$
$=$ $\zeta_{1}n^{-1/2}(\log p)^{1/2}(p^{-5/2}+p^{-1-\delta/2}).$
The second and last part in Equation 7.3:
From the proof of Lemma A3. in Cheng et al. (2023), we can conclude,
$\frac{1}{n}\sum_{i=1}^{n}v_{i}^{-1}(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{j}\bm{W}_{i})(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{l}\bm{W}_{i})^{\top}\left\{\|\mathbf{D}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|^{-3}-p^{-3/2}\right\}=O_{p}(\zeta_{1}p^{-1-\delta/2}),$
and
$n^{-1}p^{-3/2}\sum_{i=1}^{n}v_{i}^{-1}(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{j}\bm{W}_{i})(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{l}\bm{W}_{i})^{\top}\lesssim\zeta_{1}p^{-1}\left|\sigma_{j\ell}\right|+O_{p}\left(\zeta_{1}n^{-1/2}p^{-1}+\zeta_{1}p^{-7/6}\right).$
It follows that,
$\mathbf{Q}_{jl}$ $=\left(n^{-1}p^{-3/2}\sum_{i=1}^{n}v_{i}^{-1}(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{j}\bm{W}_{i})(\mathbf{D}^{-1/2}\mathbf{\Gamma}_{l}\bm{W}_{i})^{\top}+O_{p}(A_{n})\right)(1+O_{p}(n^{-1}\log p)),$
where $A_{n}=\zeta_{1}n^{-1/2}(\log p)^{1/2}(p^{-5/2}+p^{-1-\delta/2})+\zeta_{1}p^{-1-\delta/2}$.
Thus,
$\left|\mathbf{Q}_{j\ell}\right|\lesssim\zeta_{1}p^{-1}\left|\sigma_{j\ell}\right|+O_{p}\left(\zeta_{1}n^{-1/2}p^{-1}+\zeta_{1}p^{-7/6}+\zeta_{1}p^{-1-\delta/2}+\zeta_{1}n^{-1/2}(\log p)^{1/2}(p^{-5/2}+p^{-1-\delta/2})\right).$
Lemma 9.
Suppose Assumptions 1-3 holds with $a_{0}(p)\asymp p^{1-\delta}$ for some positive constant $\delta\leq 1/2$. Then, if $\log p=o(n^{1/3})$,
(i)$\|n^{-1}\sum_{i=1}^{n}\zeta_{1}^{-1}\hat{\bm{U}}_{i}\|_{\infty}=O_{p}\left\{n^{-1/2}\log^{1/2}(np)\right\}$,
(ii)$\|\zeta_{1}^{-1}n^{-1}\sum_{i=1}^{n}\delta_{1i}\hat{\bm{U}}_{i}\|_{\infty}=O_{p}(n^{-1})$.
Proof.  For any $j\in\{1,2,\cdots,p\}$,
$\hat{\bm{U}}_{ij}-{\bm{U}}_{ij}$ $=\frac{\|\mathbf{D}^{-1/2}X_{i}\|}{\|\hat{\mathbf{D}}^{-1/2}X_{i}\|}\cdot\frac{d_{j}}{\hat{d}_{j}}{\bm{U}}_{ij}-{\bm{U}}_{ij}$ (7.6)
$\leq(1+H)(1+H){\bm{U}}_{ij}-{\bm{U}}_{ij}$
$=H_{u}{\bm{U}}_{ij},$
where $H_{u}=O_{p}(H^{2}+2H)=O_{p}(n^{-1/2}(\log p)^{1/2})$. Thus, $\hat{\bm{U}}_{i}-{\bm{U}}_{i}=H_{u}{\bm{U}}_{i}$.
(i)By Equation 7.6, we have,
$\left\|n^{-1}\sum_{i=1}^{n}\zeta_{1}^{-1}\hat{\bm{U}}_{i}\right\|_{\infty}=\left\|n^{-1}\sum_{i=1}^{n}\zeta_{1}^{-1}(1+H_{u}){\bm{U}}_{i}\right\|_{\infty}$
$\leq$ $|1+H_{u}|\cdot\left\|n^{-1}\sum_{i=1}^{n}\zeta_{1}^{-1}{\bm{U}}_{i}\right\|_{\infty}=O_{p}\left\{n^{-1/2}\log^{1/2}(np)\right\}.$
(ii)Similarly,
$\left|\zeta_{1}^{-1}n^{-1}\sum_{i=1}^{n}\delta_{1i}\hat{\bm{U}}_{i}\right|_{\infty}\leq|1+H_{u}|\cdot\left\|\zeta_{1}^{-1}n^{-1}\sum_{i=1}^{n}\delta_{1i}{\bm{U}}_{i}\right\|_{\infty}$
$\leq$ $O_{p}(n^{-1}(1+n^{-1/2}\log^{1/2}p))=O_{p}(n^{-1}).$
7.1.1 Proof of Lemma 1(Bahadur representation)
Proof.  As $\bm{\theta}$ is a location parameter, we assume $\bm{\theta}=0$ without loss of generality. Then ${\bm{U}}_{i}$ can be written as ${\bm{U}}_{i}=\mathbf{D}^{-1/2}\bm{X}_{i}/\|\mathbf{D}^{-1/2}\bm{X}_{i}\|=\mathbf{D}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}/\|\mathbf{D}^{-1/2}\mathbf{\Gamma}\bm{W}_{i}\|$ for $i=1,2,\cdots,n$. The estimator $\hat{\bm{\theta}}$ satisfies $\sum_{i=1}^{n}U(\hat{\mathbf{D}}^{-1/2}(\bm{X}_{i}-\hat{\bm{\theta}}))=0$, which is is equivalent to
$\frac{1}{n}\sum_{i=1}^{n}(\hat{\bm{U}}_{i}-\hat{R}_{i}^{-1}\hat{\mathbf{D}}^{-1/2}\hat{\bm{\theta}})(1-2\hat{R}_{i}^{-1}\hat{\bm{U}}_{i}^{\top}\hat{\mathbf{D}}^{-1/2}\hat{\bm{\theta}}+\hat{R}_{i}^{-2}\hat{\bm{\theta}}^{\top}\hat{\mathbf{D}}^{-1}\hat{\bm{\theta}})^{-1/2}=0.$
From the proof of lemma A.3 in Feng et al. (2016), we can see that $\|\hat{\bm{\theta}}\|=O_{p}(\zeta_{1}^{-1}n^{-1/2})$. By the first-order Taylor expansion, the above equation can be rewritten as:
$n^{-1}\sum_{i=1}^{n}\left(\hat{\bm{U}}_{i}-\hat{R}_{i}^{-1}\hat{\mathbf{D}}^{-1/2}\hat{\bm{\theta}}\right)\left(1+\hat{R}_{i}^{-1}\hat{\bm{U}}_{i}^{\top}\hat{\mathbf{D}}^{1/2}\hat{\bm{\theta}}-2^{-1}\hat{R}_{i}^{-2}\left\|\hat{\mathbf{D}}^{-1/2}\hat{\bm{\theta}}\right\|^{2}+\delta_{1i}\right)=0,$
where $\delta_{1i}=O_{p}\left\{\left(\hat{R}_{i}^{-1}\hat{\bm{U}}_{i}^{\top}\hat{\mathbf{D}}^{1/2}\hat{\bm{\theta}}-2^{-1}\hat{R}_{i}^{-2}\left\|\hat{\mathbf{D}}^{-1/2}\hat{\bm{\theta}}\right\|^{2}\right)^{2}\right\}=O_{p}\left(n^{-1}\right)$, which implies
$\frac{1}{n}\sum_{i=1}^{n}(1-\frac{1}{2}\hat{R}_{i}^{-2}\hat{\bm{\theta}}^{\top}\hat{\mathbf{D}}^{-1}\hat{\bm{\theta}}+\delta_{1i})\hat{\bm{U}}_{i}+\frac{1}{n}\sum_{i=1}^{n}\hat{R}_{i}^{-1}(\hat{\bm{U}}_{i}^{\top}\hat{\mathbf{D}}^{-1/2}\hat{\bm{\theta}})\hat{U}_{i}$ (7.7)
$=$ $\frac{1}{n}\sum_{i=1}^{n}(1+\delta_{1i}+\delta_{2i})\hat{R}_{i}^{-1}\hat{\mathbf{D}}^{-1/2}\hat{\bm{\theta}},$
where $\delta_{2i}=O_{p}(\hat{R}_{i}^{-1}\hat{\bm{U}}_{i}^{\top}\hat{\mathbf{D}}^{1/2}\hat{\bm{\theta}}-2^{-1}\hat{R}_{i}^{-2}\left\|\hat{\mathbf{D}}^{-1/2}\hat{\bm{\theta}}\right\|^{2})=O_{p}(\delta_{1i}^{1/2})$.
Similar with Cheng et al. (2023), by Assumption 2 and Markov inequality, we have that: $\max R_{i}^{-2}=O_{p}(\zeta_{1}^{2}n^{1/2})$, $\max\delta_{1i}=O_{p}(\|\hat{\mathbf{D}}^{-1/2}\bm{\theta}\|^{2}\max\hat{R}_{i}^{-2}=O_{p}(n^{-1/2})$ and $\max\delta_{2i}=O_{p}(n^{-1/4})$.
Considering the second term in Equation 7.7,
$\frac{1}{n}\sum_{i=1}^{n}\hat{R}_{i}^{-1}(\hat{U}_{i}^{\top}\hat{\mathbf{D}}^{-1/2}\hat{\bm{\theta}})\hat{U}_{i}=\frac{1}{n}\sum_{i=1}^{n}\hat{R}_{i}^{-1}(\hat{U}_{i}\hat{U}_{i}^{\top}\hat{\mathbf{D}}^{-1/2})\hat{\bm{\theta}}=\hat{Q}\hat{\mathbf{D}}^{-1/2}\hat{\bm{\theta}},$
where $\hat{\bm{Q}}=\frac{1}{n}\sum_{i=1}^{n}\hat{R}_{i}^{-1}\hat{\bm{U}}_{i}\hat{\bm{U}}_{i}^{\top}$. From Lemma 8 we acquire,
$\left|\mathbf{Q}_{j\ell}\right|\lesssim\zeta_{1}p^{-1}\left|\omega_{j\ell}\right|+O_{p}\left(\zeta_{1}n^{-1/2}p^{-1}+\zeta_{1}p^{-7/6}+\zeta_{1}p^{-1-\delta/2}+\zeta_{1}n^{-1/2}(\log p)^{1/2}(p^{-5/2}+p^{-1-\delta/2})\right),$
and this implies that,
$\left\|\mathbf{Q}\hat{\mathbf{D}}^{-1/2}\hat{\bm{\theta}}\right\|_{\infty}$ (7.8)
$\leqslant$ $\|\mathbf{Q}\|_{1}\left\|\hat{\mathbf{D}}^{-1/2}\hat{\bm{\theta}}\right\|_{\infty}$
$\lesssim$ $\zeta_{1}p^{-1}\|\mathbf{\Omega}\|_{1}\left\|\hat{\mathbf{D}}^{-1/2}\hat{\bm{\theta}}\right\|_{\infty}$
$+$ $O_{p}\left(\zeta_{1}n^{-1/2}p^{-1}+\zeta_{1}p^{-7/6}+\zeta_{1}p^{-1-\delta/2}+\zeta_{1}n^{-1/2}(\log p)^{1/2}(p^{-5/2}+p^{-1-\delta/2})\right)\left\|\hat{\mathbf{D}}^{-1/2}\hat{\bm{\theta}}\right\|_{\infty}$
$\quad{O_{p}\left(\zeta_{1}n^{-1/2}p^{-1}+\zeta_{1}p^{-7/6}+\zeta_{1}p^{-1-\delta/2}+\zeta_{1}n^{-1/2}(\log p)^{1/2}p^{-5/2}\right)\left\|\hat{\mathbf{D}}^{-1/2}\hat{\bm{\theta}}\right\|_{\infty}}.$
According to Lemma 9, $\left\|n^{-1}\sum_{i=1}^{n}\zeta_{1}^{-1}\hat{\bm{U}}_{i}\right\|_{\infty}=O_{p}\left\{n^{-1/2}\log^{1/2}(np)\right\}$ and $\left\|\zeta_{1}^{-1}n^{-1}\sum_{i=1}^{n}\delta_{1i}\hat{\bm{U}}_{i}\right\|_{\infty}=O_{p}(n^{-1})$. In addition, we obtain,
$\left\|\zeta_{1}^{-1}n^{-1}\sum_{i=1}^{n}\hat{R}_{i}^{-2}\|\hat{\mathbf{D}}^{-1/2}\hat{\bm{\theta}}\|^{2}\hat{\bm{U}}_{i}\right\|_{\infty}\leq|1+H_{u}|\cdot\left\|\zeta_{1}^{-1}n^{-1}\sum_{i=1}^{n}\hat{R}_{i}^{-2}\|\hat{\mathbf{D}}^{-1/2}\hat{\bm{\theta}}\|^{2}{\bm{U}}_{i}\right\|_{\infty}$
$\lesssim$ $O_{p}(n^{-1})(1+{O_{p}(n^{-3/2}(\log p)^{1/2})})=O_{p}(n^{-1}).$
Considering the third term :
Using the fact that $\zeta_{1}^{-1}n^{-1}\sum_{i=1}^{n}R_{i}^{-1}=1+O_{p}(n^{-1/2})$ and Equation 7.5, we have
[…]
We final obtain:
$\left.\hat{\mathbf{D}}^{-1/2}\hat{\bm{\theta}}\right|_{\infty}$ $\lesssim\left\|\zeta_{1}^{-1}n^{-1}\sum_{i=1}^{n}\hat{\bm{U}}_{i}\right\|_{\infty}+\zeta_{1}^{-1}\left\|\mathbf{Q}\hat{\mathbf{D}}^{-1/2}\hat{\bm{\theta}}\right\|_{\infty}$
$\lesssim p^{-1}a_{0}(p)\left\|\hat{\mathbf{D}}^{-1/2}\hat{\bm{\theta}}\right\|_{\infty}+O_{p}\left\{n^{-1/2}\log^{1/2}(np)\right\}$
$\quad+O_{p}\left(n^{-1/2}+p^{-(1/6\wedge\delta/2)}+n^{-1/2}(\log p)^{1/2}p^{-3/2}\right)\left\|\hat{\mathbf{D}}^{-1/2}\hat{\bm{\theta}}\right\|_{\infty}.$
Thus we conclude that:
$\left\|\hat{\mathbf{D}}^{-1/2}\hat{\bm{\theta}}\right\|_{\infty}=O_{p}(n^{-1/2}\log^{1/2}(np)),$
as $a_{0}(p)\asymp p^{1-\delta}$.
In addition, we have
$\left\|\zeta_{1}^{-1}\mathbf{Q}\hat{\mathbf{D}}^{-1/2}\hat{\bm{\theta}}\right\|_{\infty}=O_{p}\left(n^{-1}\log^{1/2}(np)+n^{-1/2}p^{-(1/6\wedge\delta/2)}\log^{1/2}(np)+{n^{-1}(\log p)^{1/2}p^{-3/2}\log^{1/2}(np)}\right),$
and
$n^{-1}\sum_{i=1}^{n}\hat{R}_{i}^{-1}\left(1+\delta_{1i}+\delta_{2i}\right)$
$=$ $\zeta_{1}\left\{1+O_{p}\left(n^{-1/4}\right)\right\}\left\{1+O_{p}(n^{-1/2}(\log p)^{1/2}\right\}$
$=$ $\zeta_{1}\left\{1+{O_{p}(n^{-1/4}+n^{-1/2}(\log p)^{1/2})}\right\}.$
Finally, we can write
$n^{1/2}\hat{\mathbf{D}}^{-1/2}(\hat{\bm{\theta}}-\bm{\theta})=n^{-1/2}\zeta_{1}^{-1}\sum_{i=1}^{n}{\bm{U}}_{i}+C_{n},$ (7.9)
where
$\|C_{n}\|_{\infty}=$ $O_{p}(n^{-1/2}\log^{1/2}(np)+p^{-(1/6\wedge\delta/2)}\log^{1/2}(np)+n^{-1/2}(\log p)^{1/2}p^{-3/2}\log^{1/2}(np))$
$+O_{p}(n^{-1/4}\log^{1/2}(np)+n^{-1/2}(\log p)^{1/2}\log^{1/2}(np))$
$=$ $O_{p}(n^{-1/4}\log^{1/2}(np)+p^{-(1/6\wedge\delta/2)}\log^{1/2}(np)+n^{-1/2}(\log p)^{1/2}\log^{1/2}(np)).$
7.1.2 Proof of Lemma 2(Gaussian approximation)
Proof.  Let $L_{n,p}=n^{-1/4}\log^{1/2}(np)+p^{-(1/6\wedge\delta/2)}\log^{1/2}(np)+n^{-1/2}(\log p)^{1/2}\log^{1/2}(np)$. Then for any sequence $\eta_{n}\rightarrow\infty$ and any $t\in\mathbb{R}^{p}$,
$\mathbb{P}(n^{1/2}\hat{\mathbf{D}}^{-1/2}(\hat{\bm{\theta}}-\bm{\theta})\leq t)$ $=\mathbb{P}(n^{-1/2}\zeta_{1}^{-1}\sum_{i=1}^{n}{\bm{U}}_{i}+C_{n}\leq t)$
$\leq\mathbb{P}(n^{-1/2}\zeta_{1}^{-1}\sum_{i=1}^{n}{\bm{U}}_{i}\leq t+\eta_{n}L_{n,p})+\mathbb{P}(\|C_{n}\|_{\infty}>\eta_{n}L_{n,p}).$
According to Lemma 11, $\mathbb{E}\{(\zeta_{1}^{-1}{\bm{U}}_{i,j})^{4}\}\lesssim\bar{M}^{2}$ and $\mathbb{E}\{(\zeta_{1}^{-1}{\bm{U}}_{i,j})^{2}\}\gtrsim\underline{m}$ for all $i=1,2,\cdots,n$, $j=1,2,\cdots,p$, and the Gaussian approximation for independent partial sums in Koike (2021), let $\bm{G}\sim N\left(0,\zeta_{1}^{-2}\mathbf{\Sigma}_{u}\right)$ with $\mathbf{\Sigma}_{u}=\mathbb{E}\left({\bm{U}}_{1}{\bm{U}}_{1}^{\top}\right)$, we have
$\mathbb{P}(n^{1/2}\zeta_{1}^{-1}\sum_{i=1}^{n}{\bm{U}}_{i}\leq t+\eta_{n}L_{n,p})$ $\leq\mathbb{P}(\bm{G}\leq t+\eta_{n}L_{n,p})+O(\{n^{-1}\log^{5}(np)\}^{1/6})$
$\leq\mathbb{P}(\bm{G}\leq t)+O\{\eta_{n}L_{n,p}\log^{1/2}(p)\}+O(\{n^{-1}\log^{5}(np)\}^{1/6})$
where the second inequality holds from Nazarov’s inequality in Lemma 12. Thus,
$\mathbb{P}(n^{1/2}\hat{\mathbf{D}}^{-1/2}(\hat{\bm{\theta}}-\bm{\theta})\leq t)\leq$ $\mathbb{P}(\bm{G}\leq t)+O\{\eta_{n}L_{n,p}\log^{1/2}(p)\}+O(\{n^{-1}\log^{5}(np)\}^{1/6})$
$+\mathbb{P}(|C_{n}|_{\infty}>\eta_{n}l_{n,p}).$
On the other hand, we have
$\mathbb{P}(n^{1/2}\hat{\mathbf{D}}^{-1/2}(\hat{\bm{\theta}}-\bm{\theta})\leq t)\geq$ $\mathbb{P}(\bm{G}\leq t)-O\{\eta_{n}L_{n,p}\log^{1/2}(p)\}-O(\{n^{-1}\log^{5}(np)\}^{1/6})$
$-\mathbb{P}(\|C_{n}\|_{\infty}>\eta_{n}l_{n,p}).$
where $\mathbb{P}(\|C_{n}\|_{\infty}>\eta_{n}l_{n,p})\rightarrow 0$ as $n\rightarrow\infty$ by Lemma 1.
Then we have that, if $\log p=o(n^{1/5})$ and $\log n=o(p^{1/3\wedge\delta})$,
$\sup_{t\in\mathbb{R}^{p}}|\mathbb{P}(n^{1/2}\hat{\mathbf{D}}^{-1/2}(\hat{\bm{\theta}}-\bm{\theta})\leq t)-\mathbb{P}(\bm{G}\leq t)|\rightarrow 0.$
Further,
$\rho_{n}(\mathcal{A}^{re})=\sup_{A\in\mathcal{A}^{re}}|\mathbb{P}(n^{1/2}\hat{\mathbf{D}}^{-1/2}(\hat{\bm{\theta}}-\bm{\theta})\in A)-\mathbb{P}(\bm{G}\in A)|\rightarrow 0,$
by the Corollary 5.1 in Chernozhukov et al. (2017).
7.1.3 Proof of Lemma 3(Variance approximation)
Proof.$\mathbb{E}Z_{j}^{2}=\zeta_{1}^{-2}E(R_{i}^{2})^{-1}\leq\bar{B}$ by Assumption 2 and $\mathbb{E}[\max_{1\leq j\leq p}Z_{j}]\asymp(\sqrt{\log p+\log\log p})$ by Theorem 2 in Feng et al. (2022a). Let $\Delta_{0}=\max_{1\leq j,k\leq p}|p(\mathbb{E}U_{1}U_{1}^{\top})_{j,k}-R_{j,k}|$, by Lemma 11,
$\Delta_{0}=$ $\max_{1\leq j,k\leq p}|p(\mathbb{E}{\bm{U}}_{1}{\bm{U}}_{1}^{\top})_{j,k}-\mathbf{R}_{j,k}|=O(p^{-\delta/2}).$
According to Lemma 13 ,we get
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(\|\bm{Z}\|_{\infty}\leqslant t\right)-\mathbb{P}\left(\|\bm{G}\|_{\infty}\leqslant t\right)\right|\leqslant C^{\prime}n^{-1/3}\left(1\vee\log\left(np\right)\right)^{2/3}\rightarrow 0.$
7.2 Proof of main results
7.2.1 Proof of Theorem 1(Limit distribution of maxima)
Proof.
$\widetilde{\rho}_{n}$ $=\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(n^{1/2}\left\|\hat{\mathbf{D}}^{-1/2}(\hat{\bm{\theta}}-\bm{\theta})\right\|_{\infty}\leqslant t\right)-\mathbb{P}\left(\|\bm{Z}\|_{\infty}\leqslant t\right)\right|$
$\leq\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(n^{1/2}\left\|\hat{\mathbf{D}}^{-1/2}(\hat{\bm{\theta}}-\bm{\theta})\right\|_{\infty}\leqslant t\right)-\mathbb{P}\left(\|\bm{G}\|_{\infty}\leqslant t\right)\right|+\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(\left\|\bm{G}\right\|_{\infty}\leqslant t\right)-\mathbb{P}\left(\|\bm{Z}\|_{\infty}\leqslant t\right)\right|$
$\rightarrow 0.$
The last step holds from Lemma 2 and 3.
7.2.2 Proof of Theorem 2(Exact limit distribution of maxima)
Proof.  According to the Theorem 2 in Feng et al. (2022a), we have
$\mathbb{P}(p\zeta_{1}^{2}\max_{1\leq i\leq p}Z_{i}^{2}-2\log p+\log\log p\leq x)\rightarrow F(x)=\exp\left\{-\frac{1}{\sqrt{\pi}}e^{-x/2}\right\},$
a cdf of the Gumbel distribution, as $p\rightarrow\infty$. Thus, according to Lemma 7 and Theorem 1.
$\left|\mathbb{P}(T_{MAX}-2\log p+\log\log p\leq x)-F(x)\right|$
$\leq$ $\left|\mathbb{P}(\zeta_{1}^{2}\hat{\zeta}_{1}^{-2}T_{MAX}-2\log p+\log\log p\leq x)-F(x)\right|+o(1)$
$\leq$ $\left|\mathbb{P}(\zeta_{1}^{2}\hat{\zeta}_{1}^{-2}T_{MAX}-2\log p+\log\log p\leq x)-\mathbb{P}(p\zeta_{1}^{2}\max_{1\leq i\leq p}Z_{i}^{2}-2\log p+\log\log p\leq x)\right|$
$+\left|\mathbb{P}(p\zeta_{1}^{2}\max_{1\leq i\leq p}Z_{i}^{2}-2\log p+\log\log p\leq x)-F(x)\right|+o(1)\rightarrow 0,$
for any $x\in\mathbb{R}$.
7.2.3 Proof of Theorem 3(Consistency for max-type test)
Proof.  Recall that $u_{p}(y)=y+2\log p-\log\log p$, $T_{MAX}=n\|\hat{\mathbf{D}}^{-1/2}\hat{\bm{\theta}}\|_{\infty}^{2}\hat{\zeta}_{1}^{2}p\cdot\left(1-n^{-1/2}\right)$. In order to make the following proof process briefly, we abbreviate $u_{p}(0)$ to $u_{p}$, define $\widetilde{q}_{1-\alpha}=(\max\{q_{1-\alpha}+u_{p},0\})^{1/2}=O_{p}(\{\log p-2\log\log(1-\alpha)^{-1}\}^{1/2})$, $T=T_{MAX}^{1/2}=n^{1/2}\|\hat{\mathbf{D}}^{-1/2}\hat{\bm{\theta}}\|_{\infty}\cdot\hat{\zeta}_{1}p^{1/2}(1-n^{-1/2})^{1/2}$ and $T^{c}=n^{1/2}\|\hat{\mathbf{D}}^{-1/2}(\hat{\bm{\theta}}-\bm{\theta})\|_{\infty}\cdot\hat{\zeta}_{1}p^{1/2}(1-n^{-1/2})^{1/2}$, which has the same distribution of $T$ under $H_{0}$.
It is clear that, $T\geq n^{1/2}\|\hat{\mathbf{D}}^{-1/2}\bm{\theta}\|_{\infty}\cdot\hat{\zeta}_{1}p^{1/2}(1-n^{-1/2})^{1/2}-T^{c}$. Combined with Assumption 2 and Lemma 6, we get
$\mathbb{P}(T_{MAX}-u_{p}\geq q_{1-\alpha}\mid H_{1})$
$\geq$ $\mathbb{P}(n^{1/2}\|\hat{\mathbf{D}}^{-1/2}\bm{\theta}\|_{\infty}\cdot\hat{\zeta}_{1}p^{1/2}(1-n^{-1/2})^{1/2}-T^{c}\geq\widetilde{q}_{1-\alpha}\mid H_{1})$
$=$ $\mathbb{P}(T^{c}\leq n^{1/2}\|\hat{\mathbf{D}}^{-1/2}\bm{\theta}\|_{\infty}\cdot\hat{\zeta}_{1}p^{1/2}(1-n^{-1/2})^{1/2}-\widetilde{q}_{1-\alpha}\mid H_{1})$
$\geq$ $\mathbb{P}(T^{c}\leq n^{1/2}\left(\|{\mathbf{D}}^{-1/2}\bm{\theta}\|_{\infty}-\|(\hat{\mathbf{D}}^{-1/2}-{\mathbf{D}}^{-1/2})\bm{\theta}\|_{\infty}\right)\cdot\hat{\zeta}_{1}p^{1/2}(1-n^{-1/2})^{1/2}-\widetilde{q}_{1-\alpha}\mid H_{1})$
$\geq$ $\mathbb{P}(T^{c}\leq n^{1/2}\|{\mathbf{D}}^{-1/2}\bm{\theta}\|_{\infty}\cdot(1+O_{p}(n^{1/2}\log^{1/2}(np)))\cdot\hat{\zeta}_{1}p^{1/2}(1-n^{-1/2})^{1/2}-\widetilde{q}_{1-\alpha}\mid H_{1})\rightarrow 1,$
if $\|\bm{\theta}\|_{\infty}\geq\widetilde{C}n^{-1/2}\{\log p-2\log\log(1-\alpha)^{-1}\}^{1/2}$ for some large enough constant $\widetilde{C}$.
The last inequality holds since
$\|(\hat{\mathbf{D}}^{-1/2}-{\mathbf{D}}^{-1/2})\bm{\theta}\|_{\infty}$ $=\max_{i=1,2,\cdots,p}\frac{\hat{d}_{i}-d_{i}}{\hat{d}_{i}d_{i}}\theta_{i}\leq\max_{i=1,2,\cdots,p}|1-\frac{d_{i}}{\hat{d}_{i}}|\cdot\|{\mathbf{D}}^{-1/2}\bm{\theta}\|_{\infty}$
$\leq O_{p}(n^{-1/2}\log^{1/2}(np))\|{\mathbf{D}}^{-1/2}\bm{\theta}\|_{\infty}.$
7.2.4 Proof of Theorem 5(Without elliptical distribution)
Theorem 4 is the special case of Theorem 5 with $\bm{\theta}=0$, so we only need to show that Theorem 5 holds.
Proof.  The following proof is based on the idea of the proof in article Feng and Sun (2016), with modifications in some equations. We restate the equations in Feng and Sun (2016) on $U(\hat{\mathbf{D}}_{ij}^{-1/2}\bm{X}_{i})^{\top}U(\hat{\mathbf{D}}_{ij}^{-1/2}\bm{X}_{i})$. By the definition, we have
$\frac{2}{n(n-1)}\sum_{1\leq i<j\leq n}U\left(\hat{\mathbf{D}}_{ij}^{-1/2}\bm{X}_{i}\right)^{\top}U\left(\hat{\mathbf{D}}_{ij}^{-1/2}\bm{X}_{j}\right)$
$=$ $\frac{2}{n(n-1)}\sum_{1\leq i<j\leq n}\left(\bm{U}_{i}+R_{i}^{-1}\mathbf{D}^{-1/2}\bm{\theta}+\left(\hat{\mathbf{D}}_{ij}^{-1/2}\mathbf{D}^{1/2}-\mathbf{I}_{p}\right)\bm{U}_{i}\right)^{\top}$
×$\times\left(\bm{U}_{j}+R_{j}^{-1}\mathbf{D}^{-1/2}\bm{\theta}+\left(\hat{\mathbf{D}}_{ij}^{-1/2}\mathbf{D}^{1/2}-\mathbf{I}_{p}\right)\bm{U}_{j}\right)\left(1+\alpha_{ij}\right)^{-1/2}\left(1+\alpha_{ji}\right)^{-1/2}$
$=$ $\frac{2}{n(n-1)}\sum_{1\leq i<j\leq n}\bm{U}_{i}^{\top}\bm{U}_{j}+\frac{2}{n(n-1)}\sum_{1\leq i<j\leq n}R_{i}^{-1}R_{j}^{-1}\bm{\theta}^{\top}\mathbf{D}^{-1}\bm{\theta}$
$+\frac{2}{n(n-1)}\sum_{1\leq i<j\leq n}\bm{U}_{i}^{\top}\bm{U}_{j}\left[\left(1+\alpha_{ij}\right)^{-1/2}\left(1+\alpha_{ji}\right)^{-1/2}-1\right]$
$+\frac{4}{n(n-1)}\sum_{1\leq i<j\leq n}\bm{U}_{i}^{\top}\left(\hat{\mathbf{D}}_{ij}^{-1/2}\mathbf{D}^{1/2}-\mathbf{I}_{p}\right)\bm{U}_{j}\left(1+\alpha_{ij}\right)^{-1/2}\left(1+\alpha_{ji}\right)^{-1/2}$
$+\frac{2}{n(n-1)}\sum_{1\leq i<j\leq n}\bm{U}_{i}^{\top}\left(\hat{\mathbf{D}}_{ij}^{-1/2}\mathbf{D}^{1/2}-\mathbf{I}_{p}\right)^{2}\bm{U}_{j}\left(1+\alpha_{ij}\right)^{-1/2}\left(1+\alpha_{ji}\right)^{-1/2}$
$+\frac{2}{n(n-1)}\sum_{1\leq i<j\leq n}R_{j}^{-1}\bm{\theta}^{\top}\mathbf{D}^{-1}\bm{\theta}\left[\left(1+\alpha_{ij}\right)^{-1/2}\left(1+\alpha_{ji}\right)^{-1/2}-1\right]$
$+\frac{4}{n(n-1)}\sum_{1\leq i<j\leq n}\bm{U}_{i}^{\top}\mathbf{D}^{-1/2}\bm{\theta}\left(1+\alpha_{ij}\right)^{-1/2}\left(1+\alpha_{ji}\right)^{-1/2}$
$+\frac{2}{n(n-1)}\sum_{1\leq i<j\leq n}\bm{U}_{i}^{\top}\left(\hat{\mathbf{D}}_{ij}^{-1/2}\mathbf{D}^{1/2}-\mathbf{I}_{p}\right)\mathbf{D}^{-1/2}\bm{\theta}\left(1+\alpha_{ij}\right)^{-1/2}\left(1+\alpha_{ji}\right)^{-1/2}$
::=:= $\frac{2}{n(n-1)}\sum_{1\leq i<j\leq n}\bm{U}_{i}^{\top}\bm{U}_{j}+\frac{2}{n(n-1)}\sum_{1\leq i<j\leq n}R_{i}^{-1}R_{j}^{-1}\bm{\theta}^{\top}\mathbf{D}^{-1}\bm{\theta}$
$+A_{n1}+A_{n2}+A_{n3}+A_{n4}+A_{n5}+A_{n6}.$
where
$\alpha_{ij}=2R_{i}^{-1}\bm{U}_{i}^{\top}\left(\hat{\mathbf{D}}_{ij}^{-1/2}-\mathbf{D}^{-1/2}\right)\bm{X}_{i}+R_{i}^{-2}\left\|\left(\hat{\mathbf{D}}_{ij}^{-1/2}-\mathbf{D}^{-1/2}\right)\bm{X}_{i}\right\|^{2}+\left.2R_{i}^{-1}\mathbf{D}^{-1/2}\bm{\theta}+R_{i}^{-2}\bm{\theta}^{\top}\mathbf{D}^{-1}\bm{\theta}\right)^{-1/2}.$
Note that $R_{i}^{-1}\bm{U}_{i}^{\top}\left(\hat{\mathbf{D}}_{ij}^{-1/2}-\mathbf{D}^{-1/2}\right)\bm{X}_{i}=$ $\bm{U}_{i}^{\top}\left(\hat{\mathbf{D}}_{ij}^{-1/2}\mathbf{D}^{1/2}-\mathbf{I}_{p}\right)\bm{U}_{i}+R_{i}^{-1}\bm{U}_{i}^{\top}\left(\hat{\mathbf{D}}_{ij}^{-1/2}-\mathbf{D}^{-1/2}\right)\bm{\theta}=O_{p}\left(n^{-1/2}\left(\log p\right)^{1/2}\right)$ and $R_{i}^{-2}\left\|\left(\hat{\mathbf{D}}_{ij}^{-1/2}-\mathbf{D}^{-1/2}\right)\bm{X}_{i}\right\|^{2}=O_{p}\left(n^{-1}\log p\right)$ by Lemma 6. By Assumption 7 and Equation 3.1, $R_{i}^{-1}\mathbf{D}^{-1/2}\bm{\theta}=O_{p}\left(\sigma_{n}^{1/2}\right)=O_{p}\left(n^{-1}\right)$ and $R_{i}^{-2}\bm{\theta}^{\top}\mathbf{D}^{-1}\bm{\theta}=$ $O_{p}\left(\sigma_{n}\right)=O_{p}\left(n^{-2}\right)$ where $\sigma_{n}^{2}=\frac{2}{n(n-1)p^{2}}\operatorname{tr}\left(\mathbf{R}^{2}\right)$. So $\alpha_{ij}=O_{p}\left(n^{-1/2}\left(\log p\right)^{1/2}\right)$.
Similarly, we will show that $A_{n1}=o_{p}\left(\sigma_{n}\right)$. Under some calculations, we get $\mathbb{E}[(\bm{U}_{i}^{\top}\bm{U}_{j})^{2}]=\text{tr}(\mathbf{\Sigma}^{2}_{u})$. By Lemma 11, we find that $\mathbf{\Sigma}_{u,i,j}=p^{-1}\sigma_{i,j}+O(p^{-1-\delta/2})$. Thus we have,
$\text{tr}(\mathbf{\Sigma}^{2}_{u})$ $=\sum_{i=1}^{p}\sum_{j=1}^{p}\mathbf{\Sigma}_{u,i,j}^{2}=\sum_{i=1}^{p}\sum_{j=1}^{p}\left(p^{-2}\sigma^{2}_{i,j}+\sigma_{i,j}O(p^{-2-\delta/2})\right)$ (7.10)
$=\sum_{i=1}^{p}\sum_{j=1}^{p}p^{-2}\sigma^{2}_{i,j}+\sum_{p^{-\delta/2}=O(\sigma_{ij})}\sigma_{i,j}O(p^{-2-\delta/2})$
$+\sum_{\sigma_{ij}\in[C_{1}\frac{O(\text{tr}(\mathbf{R}^{2}))}{p^{2-\delta/2}},C_{2}p^{-\delta/2}]}\sigma_{i,j}O(p^{-2-\delta/2})+\sum_{\sigma_{ij}=O(\frac{O(\text{tr}(\mathbf{R}^{2}))}{p^{2-\delta/2}})}\sigma_{i,j}O(p^{-2-\delta/2})$
$=p^{-2}\text{tr}(\mathbf{R}^{2})(1+O(1))+O(p^{-2-\delta/2})\cdot\frac{p^{2-\delta/2}}{O(\text{tr}(\mathbf{R}^{2}))}\cdot o(\frac{p^{2}}{n})+p^{-2}O(\text{tr}(\mathbf{R}^{2}))$
$=O(p^{-2}\text{tr}(\mathbf{R}^{2})).$
By the Cauchy inequality,
$E\left(A_{n1}^{2}\right)$ $=O\left(n^{-4}\right)\sum_{i<j}E\left\{\bm{U}_{i}^{T}\bm{U}_{j}\left[\left(1+\alpha_{ij}\right)^{-1/2}\left(1+\alpha_{ji}\right)^{-1/2}-1\right]\right\}^{2}$
$\leq O\left(n^{-2}\right)E\left(\bm{U}_{i}^{T}\bm{U}_{j}\right)^{2}E\left[\left(1+\alpha_{ij}\right)^{-1/2}\left(1+\alpha_{ji}\right)^{-1/2}-1\right]^{2}$
$=O\left(n^{-3}\log p\left(p^{-2}\text{tr}(\mathbf{R}^{2})+O(p^{-2-\delta/2})(p+po(\frac{p}{n^{1/2}}))\right)\right)=o\left(\sigma_{n}^{2}\right).$
$A_{n2}$ $=\frac{4}{n(n-1)}\sum_{i<j}\bm{U}_{i}^{\top}\left(\hat{\mathbf{D}}_{ij}^{-1/2}\mathbf{D}^{1/2}-\mathbf{I}_{p}\right)\bm{U}_{j}$
$\quad+\frac{4}{n(n-1)}\sum_{i<j}\bm{U}_{i}^{\top}\left(\hat{\mathbf{D}}_{ij}^{-1/2}\mathbf{D}^{1/2}-\mathbf{I}_{p}\right)\bm{U}_{j}\left[\left(1+\alpha_{ij}\right)^{-1/2}\left(1+\alpha_{ji}\right)^{-1/2}-1\right]$
$:=G_{n1}+G_{n2}.$
By Lemma 6 and Equation 7.10, $\mathbb{E}((\bm{U}_{i}^{\top}\left(\hat{\mathbf{D}}_{ij}^{-1/2}\mathbf{D}^{1/2}-\mathbf{I}_{p}\right)\bm{U}_{j})^{2})\leq O(n^{-1}\log p\text{ tr}(\mathbf{\Sigma}_{u}^{2}))=o(p^{-2}\text{tr}(\mathbf{R}^{2}))$. Then we obtain $G_{n1}=o_{p}(\sigma_{n})$. Similar to $A_{n1}$, we can show $G_{n2}=o_{p}(\sigma_{n})$. Taking the same procedure as $A_{n2}$, we can obtain $A_{n3}=o_{p}(\sigma_{n})$. Similarly to the processing of Equation 7.1, we get
$\frac{2}{n(n-1)}\sum_{1\leq i<j\leq n}\bm{U}_{i}^{\top}\bm{U}_{j}=\frac{2}{n(n-1)}\sum_{1\leq i<j\leq n}(\mathbf{D}^{-1/2}\mathbf{\Gamma}U(\bm{W}_{i}))^{\top}\mathbf{D}^{-1/2}\mathbf{\Gamma}U(\bm{W}_{j})+o_{p}(\sigma_{n}).$
We replace the Lemma 1 in Feng and Sun (2016) by Lemma 11, and final acquire
[…]
7.2.5 Proof of Theorem 6(Asymptotically independent under $H_{0}$)
Proof.  To prove $T_{SUM}$ and $T_{MAX}$ are asymptotically independent, it suffices to show that: Under $H_{0}$,
$\mathbb{P}(\frac{T_{SUM}}{\sigma_{n}}\leq x,T_{MAX}-2\log p+\log\log p\leq y)\\ \rightarrow\Phi(x)\cdot\exp\left\{-\frac{1}{\sqrt{\pi}}e^{-y/2}\right\}.$ (7.11)
Let $u_{p}(y)=y+2\log p-\log\log p$, and we rewrite Equation 7.11 as
$\mathbb{P}(\frac{T_{SUM}}{\sigma_{n}}\leq x,T_{MAX}\leq u_{p}(y))\rightarrow\Phi(x)\cdot\exp\left\{-\frac{1}{\sqrt{\pi}}e^{-y/2}\right\}.$ (7.12)
From the proof of Theorem 2 in Feng and Sun (2016), we acquire
$T_{SUM}=\frac{2}{n(n-1)}\sum\sum_{i<j}\bm{U}_{i}^{\top}\bm{U}_{j}+o_{p}(\sigma_{n}),$ (7.13)
and it’s easy to find that $\sigma_{n}^{2}=\frac{2}{n(n-1)p}+o(\frac{1}{n^{3}})$ according to Assumption 7. Combined with Lemma 1, it suffice to show,
$\mathbb{P}(\frac{\frac{2}{n(n-1)}\sum\sum_{i<j}\bm{U}_{i}^{\top}\bm{U}_{j}}{\sigma_{n}}+o_{p}(1)\leq x$ (7.14)
$\qquad,p\left\|n^{-1/2}\sum_{i=1}^{n}\bm{U}_{i}\right\|_{\infty}^{2}+O_{p}(L_{n,p})\leq u_{p}(y))$
$\rightarrow\Phi(x)\cdot\exp\left\{-\frac{1}{\sqrt{\pi}}e^{-y/2}\right\}.$
We next prove that,
$\mathbb{P}(\sqrt{\frac{n}{n-1}}\left(\frac{\|\sqrt{\frac{p}{n}}\sum_{i=1}^{n}\bm{U}_{i}\|_{2}^{2}-p}{\sqrt{2\text{tr}(\mathbf{R}^{2})}}\right)\leq x,\left\|\sqrt{\frac{p}{n}}\sum_{i=1}^{n}\bm{U}_{i}\right\|_{\infty}^{2}\leq u_{p}(y))$ (7.15)
$\rightarrow\Phi(x)\cdot\exp\left\{-\frac{1}{\sqrt{\pi}}e^{-y/2}\right\}.$
When Equation 7.15 holds, combined with $O_{p}(L_{n,p})=o_{p}(1)$, Equation 7.14 holds obviously, which means that the independence of $T_{SUM}$ and $T_{MAX}$ follows.
Proof of Equation 7.15: From the Theorem 2 in Feng et al. (2022a), the Equation 7.15 holds if $\bm{U}_{i}$ follows the normal distribution. We then investigate the non-normal case. Let $\bm{\xi}_{i}=\bm{U}_{i}\in R^{p},i=1,2,\cdots,n$. For $\bm{z}=(z_{1},\cdots,z_{q})^{\top}\in R^{q}$, we consider a smooth approximation of the maximum function, namely,
$F_{\beta}(\bm{z}):=\beta^{-1}\log(\sum_{j=1}^{q}\exp(\beta z_{j})),$
where $\beta>0$ is the smoothing parameter that controls the level of approximation. An elementary calculation shows that for all $z\in R^{q}$,
$0\leq F_{\beta}(\bm{z})-\max_{1\leq j\leq q}z_{j}\leq\beta^{-1}\log q.$
Define $\sigma_{S}^{2}=2n^{2}\text{ tr }(R^{2})$,
$W(\bm{x}_{1},\cdots,\bm{x}_{n})$ $=\frac{\|\sqrt{\frac{p}{n}}\sum_{i=1}^{n}\bm{x}_{i}\|_{2}^{2}-p}{\sqrt{2\text{tr}(\mathbf{R}^{2})}}$
=$\frac{p\sum_{i\not=j}\bm{x}_{i}^{\top}\bm{x}_{j}}{\sqrt{2n^{2}\text{ tr }(\mathbf{R}^{2})}}:=\frac{p\sum_{i\not=j}\bm{x}_{i}^{\top}\bm{x}_{j}}{\sigma_{S}},$
$V(\bm{x}_{1},\cdots,\bm{x}_{n})$ $=\beta^{-1}\log(\sum_{j=1}^{p}\exp(\beta\sqrt{\frac{p}{n}}\sum_{i=1}^{n}\bm{x}_{i,j})).$
By setting $\beta=n^{1/8}\log(n)$, Equation 7.15 is equivalent to
$P(W(\bm{\xi}_{1},\cdots,\bm{\xi}_{p})\leq x,V(\bm{\xi}_{1},\cdots,\bm{\xi}_{p})\leq u_{p}(y))\rightarrow\Phi(x)\cdot\exp(-\exp(y)).$ (7.16)
Suppose $\{\bm{Y}_{1},\bm{Y}_{2},\cdots,\bm{Y}_{n}\}$ are sample from $N(0,\mathbb{E}\bm{U}_{1}^{\top}\bm{U}_{1})$ , and independent with $\bm{U}_{1},\cdots,\bm{U}_{n}$ or write as $\bm{\xi}_{1},\cdots,\bm{\xi}_{n})$. The key idea is to show that: $(W(\bm{\xi}_{1},\cdots,\bm{\xi}_{n}),V(\bm{\xi}_{1},\cdots,\bm{\xi}_{n}))$ has the same limiting distribution as $W(\bm{Y}_{1},\cdots,\bm{Y}_{n}),V(\bm{Y}_{1},\cdots,\bm{Y}_{n}))$.
Let $l^{2}_{b}(\mathbb{R})$ denote the class of bounded functions with bounded and continuous derivatives up to order 3.It is known that a sequence of randon variables $\{Z_{n}\}_{n=1}^{\infty}$ converges weakly to a random variable $Z$ if and only if for every $f\in l^{3}_{b}(\mathbb{R})$, $\mathbb{E}(f(Z_{n}))\rightarrow\mathbb{E}(f(Z))$.
It suffices to show that:
$\mathbb{E}\{f(W(\bm{\xi}_{1},\cdots,\bm{\xi}_{n}),V(\bm{\xi}_{1},\cdots,\bm{\xi}_{n}))\}-\mathbb{E}\{f(W(\bm{Y}_{1},\cdots,\bm{Y}_{n}),V(\bm{Y}_{1},\cdots,\bm{Y}_{n}))\}\rightarrow 0,$
for every $f\in l_{b}^{3}(\mathbb{R}^{2})$ as $(n,p)\rightarrow\infty$.
We introduce $\widetilde{W}_{d}=W(\bm{\xi}_{1},\cdots,\bm{\xi}_{d-1},\bm{Y}_{d},\cdots,\bm{Y}_{n})$ and $\widetilde{V}_{d}=V(\bm{\xi}_{1},\cdots,\bm{\xi}_{d-1},\bm{Y}_{d},\cdots,\bm{Y}_{n})$ for $d=1,\cdots,n+1$, $\mathcal{F}_{d}=\sigma\{\bm{\xi}_{1},\cdots,\bm{\xi}_{d-1},\bm{Y}_{d+1},\cdots,\bm{Y}_{n}\}$ for $d=1,\cdots,n$. If there is no danger of confusion, we simply write $\widetilde{W}_{d}$ and $\widetilde{V}_{d}$ as $W_{d}$ and $V_{d}$ respectively (only for this part). Then,
$\mid\mathbb{E}\{f(W(\bm{\xi}_{1},\cdots,\bm{\xi}_{n}),V(\bm{\xi}_{1},\cdots,\bm{\xi}_{n}))\}-\mathbb{E}\{f(W(\bm{Y}_{1},\cdots,\bm{Y}_{n}),V(\bm{Y}_{1},\cdots,\bm{Y}_{n}))\}\mid$
$\leq$ $\sum_{d=1}^{n}\mid\mathbb{E}\{f(W_{d},V_{d})-\mathbb{E}\{f(W_{d+1},V_{d+1})\}\mid.$
Let
$W_{d,0}$ $=\frac{2p\sum_{i<j}^{d-1}\bm{\xi}_{i}^{\top}\bm{\xi}_{j}+2p\sum_{d+1\leq i<j\leq n}\bm{Y}_{i}^{\top}\bm{Y}_{j}+2p\sum_{i=1}^{d-1}\sum_{j=d+1}^{n}\bm{\xi}_{i}^{\top}\bm{Y}_{j}}{\sigma_{S}}\in\mathcal{F}_{d},$
$V_{d,0}$ $=\beta^{-1}\log(\sum_{j=1}^{p}\exp(\beta\sqrt{\frac{p}{n}}\sum_{i=1}^{d-1}\xi_{i,j}+\beta\sqrt{\frac{p}{n}}\sum_{i=d+1}^{n}Y_{i,j}))\in\mathcal{F}_{d}.$
By Taylor expansion, we have,
$f\left(W_{d},V_{d}\right)-f\left(W_{d,0},V_{d,0}\right)=$ $f_{1}\left(W_{d,0},V_{d,0}\right)\left(W_{d}-W_{d,0}\right)+f_{2}\left(W_{d,0},V_{d,0}\right)\left(V_{d}-V_{d,0}\right)$
$+\frac{1}{2}f_{11}\left(W_{d,0},V_{d,0}\right)\left(W_{d}-W_{d,0}\right)^{2}+\frac{1}{2}f_{22}\left(W_{d,0},V_{d,0}\right)\left(V_{d}-V_{d,0}\right)^{2}$
$+\frac{1}{2}f_{12}\left(W_{d,0},V_{d,0}\right)\left(W_{d}-W_{d,0}\right)\left(V_{d}-V_{d,0}\right)$
$+O\left(\left|\left(V_{d}-V_{d,0}\right)\right|^{3}\right)+O\left(\left|\left(W_{d}-W_{d,0}\right)\right|^{3}\right),$
and
$f\left(W_{d+1},V_{d+1}\right)-f\left(W_{d,0},V_{d,0}\right)=$ $f_{1}\left(W_{d,0},V_{d,0}\right)\left(W_{d+1}-W_{d,0}\right)+f_{2}\left(W_{d,0},V_{d,0}\right)\left(V_{d+1}-V_{d,0}\right)$
$+\frac{1}{2}f_{11}\left(W_{d,0},V_{d,0}\right)\left(W_{d+1}-W_{d,0}\right)^{2}+\frac{1}{2}f_{22}\left(W_{d,0},V_{d,0}\right)\left(V_{d+1}-V_{d,0}\right)^{2}$
$+\frac{1}{2}f_{12}\left(W_{d,0},V_{d,0}\right)\left(W_{d+1}-W_{d,0}\right)\left(V_{d+1}-V_{d,0}\right)$
$+O\left(\left|\left(V_{d+1}-V_{d,0}\right)\right|^{3}\right)+O\left(\left|\left(W_{d+1}-W_{d,0}\right)\right|^{3}\right),$
where for $f:=f(x,y),f_{1}(x,y)=\frac{\partial f}{\partial x},f_{2}(x,y)=\frac{\partial f}{\partial y},f_{11}(x,y)=\frac{\partial f^{2}}{\partial^{2}x},f_{22}(x,y)=\frac{\partial f^{2}}{\partial^{2}y}$ and $f_{12}(x,y)=\frac{\partial f^{2}}{\partial x\partial y}$.
We first consider $W_{d},W_{d+1},W_{d,0}$ and notice that,
$W_{d}-W_{d,0}$ $=\frac{p\sum_{i=1}^{d-1}\bm{\xi}_{i}^{\top}\bm{Y}_{d}+p\sum_{i=d+1}^{n}\bm{Y}_{i}^{\top}\bm{Y}_{d}}{\sigma_{S}},$
$W_{d+1}-W_{d,0}$ $=\frac{p\sum_{i=1}^{d-1}\bm{\xi}_{i}^{\top}\bm{\xi}_{d}+p\sum_{i=d+1}^{n}\bm{Y}_{i}^{\top}\bm{\xi}_{d}}{\sigma_{S}}.$
Due to $\mathbb{E}\left(\bm{\xi}_{t}\right)=\mathbb{E}\left(\bm{Y}_{t}\right)=0$ and $\mathbb{E}\left(\bm{\xi}_{t}\bm{\xi}_{t}^{\top}\right)=\mathbb{E}\left(\bm{Y}_{t}\bm{Y}_{t}^{\top}\right)$, it can be verified that,
$\mathbb{E}\left(W_{d}-W_{d,0}\mid\mathcal{F}_{d}\right)=\mathbb{E}\left(W_{d+1}-W_{d,0}\mid\mathcal{F}_{d}\right)\text{ and }\mathbb{E}\left(\left(W_{d}-W_{d,0}\right)^{2}\mid\mathcal{F}_{d}\right)=\mathbb{E}\left(\left(W_{d+1}-W_{d,0}\right)^{2}\mid\mathcal{F}_{d}\right).$
Hence,
$\mathbb{E}\left\{f_{1}\left(W_{d,0},V_{d,0}\right)\left(W_{d}-W_{d,0}\right)\right\}$ $=\mathbb{E}\left\{f_{1}\left(W_{d,0},V_{d,0}\right)\left(W_{d+1}-W_{d,0}\right)\right\}\text{ and }$
$\mathbb{E}\left\{f_{11}\left(W_{d,0},V_{d,0}\right)\left(W_{d}-W_{d,0}\right)^{2}\right\}$ $=\mathbb{E}\left\{f_{11}\left(W_{d,0},V_{d,0}\right)\left(W_{d+1}-W_{d,0}\right)^{2}\right\}.$
Next we consider $V_{d}-V_{d,0}$. Let $z_{d,0,j}=\sqrt{\frac{p}{n}}\sum_{i=1}^{d-1}\xi_{i,j}+\sqrt{\frac{p}{n}}\sum_{i=d+1}^{n}Y_{i,j},z_{d,j}=z_{d,0,j}+n^{-1/2}\sqrt{p}Y_{d,j},z_{d+1,j}=z_{d,0,j}+n^{-1/2}\sqrt{p}\xi_{d,j}$. By Taylor expansion, we have that:
$V_{d}-V_{d,0}=$ $\sum_{l=1}^{n}\partial_{l}F_{\beta}\left(\bm{z}_{d,0}\right)\left(z_{d,l}-z_{d,0,l}\right)+\frac{1}{2}\sum_{l=1}^{n}\sum_{k=1}^{n}\partial_{k}\partial_{l}F_{\beta}\left(\bm{z}_{d,0}\right)\left(z_{d,l}-z_{d,0,l}\right)\left(z_{d,k}-z_{d,0,k}\right)$ (7.17)
$+\frac{1}{6}\sum_{l=1}^{n}\sum_{k=1}^{n}\sum_{v=1}^{n}\partial_{v}\partial_{k}\partial_{l}F_{\beta}\left(\bm{z}_{d,0}+\delta\left(\bm{z}_{d}-\bm{z}_{d,0}\right)\right)\left(z_{d,l}-z_{d,0,l}\right)\left(z_{d,k}-\bm{z}_{d,0,k}\right)\left(\bm{z}_{d,v}-\bm{z}_{d,0,v}\right),$
for some $\delta\in(0,1)$. Again, due to $\mathbb{E}\left(\bm{\xi}_{t}\right)=\mathbb{E}\left(\bm{Y}_{t}\right)=0$ and $\mathbb{E}\left(\bm{\xi}_{t}\bm{\xi}_{t}^{\top}\right)=\mathbb{E}\left(\bm{Y}_{t}\bm{Y}_{t}^{\top}\right)$, we can verify that
$\mathbb{E}\left\{\left(z_{d,l}-z_{d,0,l}\right)\mid\mathcal{F}_{d}\right\}=\mathbb{E}\left\{\left(z_{d+1,l}-z_{d,0,l}\right)\mid\mathcal{F}_{d}\right\}\text{ and }\mathbb{E}\left\{\left(z_{d,l}-z_{d,0,l}\right)^{2}\mid\mathcal{F}_{d}\right\}=\mathbb{E}\left\{\left(z_{d+1,l}-z_{d,0,l}\right)^{2}\mid\mathcal{F}_{d}\right)\text{. }$
By Lemma A.2 in Chernozhukov et al. (2013), we have,
$\left|\sum_{l=1}^{n}\sum_{k=1}^{n}\sum_{v=1}^{n}\partial_{v}\partial_{k}\partial_{l}F_{\beta}\left(\bm{z}_{d,0}+\delta\left(\bm{z}_{d}-\bm{z}_{d,0}\right)\right)\right|\leq C\beta^{2},$
for some positive constant $C$.
By Lemma 11, we have that: $\left\|\zeta_{1}^{-1}U_{i,j}\right\|_{\psi_{\alpha}}\lesssim\bar{B},$ for all $i=1,\ldots,n$ and $j=1,\ldots,p$, which means $P(\mid\sqrt{p}\xi_{i,j}\mid\geq t)\leq 2\exp(-(ct\sqrt{p}/\zeta_{1})^{\alpha})\lesssim 2\exp(-(ct)^{\alpha})$, $\mathrm{P}\left(\max_{1\leq i\leq n}\left|\sqrt{p}\xi_{ij}\right|>C\log(n)\right)\rightarrow 0$ and since $\sqrt{p}Y_{tj}\sim N(0,1)$, $\mathrm{P}\left(\max_{1\leq i\leq n}\left|\sqrt{p}Y_{ij}\right|>C\log(n)\right)\rightarrow 0$. Hence,
$\left|\frac{1}{6}\sum_{l=1}^{n}\sum_{k=1}^{n}\sum_{v=1}^{n}\partial_{v}\partial_{k}\partial_{l}F_{\beta}\left(\bm{z}_{d,0}+\delta\left(\bm{z}_{d}-\bm{z}_{d,0}\right)\right)\left(\bm{z}_{d,l}-\bm{z}_{d,0,l}\right)\left(\bm{z}_{d,k}-\bm{z}_{d,0,k}\right)\left(\bm{z}_{d,v}-\bm{z}_{d,0,v}\right)\right|$
$\leq C\beta^{2}n^{-3/2}\log^{3}(n),$
$\left|\frac{1}{6}\sum_{l=1}^{n}\sum_{k=1}^{n}\sum_{v=1}^{n}\partial_{v}\partial_{k}\partial_{l}F_{\beta}\left(\bm{z}_{d+1,0}+\delta\left(\bm{z}_{d+1}-\bm{z}_{d,0}\right)\right)\left(\bm{z}_{d+1,l}-\bm{z}_{d,0,l}\right)\left(\bm{z}_{d+1,k}-\bm{z}_{d,0,k}\right)\left(\bm{z}_{d+1,v}-\bm{z}_{d,0,v}\right)\right|$
$\leq C\beta^{2}n^{-3/2}\log^{3}(n),$
holds with probability approaching one. Consequently, we have that: with probability one,
$\left|\mathrm{E}\left\{f_{2}\left(W_{d,0},V_{d,0}\right)\left(V_{d}-V_{d,0}\right)\right\}-\mathrm{E}\left\{f_{2}\left(W_{d,0},V_{d,0}\right)\left(V_{d+1}-V_{d,0}\right)\right\}\right|\leq C\beta^{2}n^{-3/2}\log^{3}(n).$
Similarly, it can be verified that,
$\left|\mathrm{E}\left\{f_{22}\left(W_{d,0},V_{d,0}\right)\left(V_{d}-V_{d,0}\right)^{2}\right\}-\mathrm{E}\left\{f_{22}\left(W_{d,0},V_{d,0}\right)\left(V_{d+1}-V_{d,0}\right)^{2}\right\}\right|\leq C\beta^{2}n^{-3/2}\log^{3}(n),$
and
$\left|\mathrm{E}\left\{f_{12}\left(W_{d,0},V_{d,0}\right)\left(W_{d}-W_{d,0}\right)\left(V_{d}-V_{d,0}\right)\right\}-\mathrm{E}\left\{f_{12}\left(W_{d,0},V_{d,0}\right)\left(W_{d+1}-W_{d,0}\right)\left(V_{d+1}-V_{d,0}\right)\right\}\right|$
$\quad\leq C\beta^{2}n^{-3/2}\log^{3}(n).$
By Equation 7.17, $\mathrm{E}\left(\left|V_{d}-V_{d,0}\right|^{3}\right)=O\left(n^{-3/2}\log^{3}(n)\right)$. For $\mathbb{E}\left(\left(W_{d}-W_{d,0}\right)^{3}\right)$, we first calculate $\mathbb{E}\left(\left(W_{d}-W_{d,0}\right)^{4}\right)$, then it’s easy to get the order for 3-order term.
$\quad\mathbb{E}\left(\left(W_{d}-W_{d,0}\right)^{4}\right)$ $=\mathbb{E}\left(\frac{p\sum_{i=1}^{d-1}\bm{\xi}_{i}^{\top}\bm{Y}_{d}+p\sum_{i=d+1}^{n}\bm{Y}_{i}^{\top}\bm{Y}_{d}}{\sigma_{S}}\right)^{4}$ (7.18)
$=\frac{p^{4}}{2n^{4}(\text{tr }(\mathbf{R}^{2}))^{2}}\mathbb{E}\left(\sum_{i=1}^{d-1}\bm{\xi}_{i}^{\top}\bm{Y}_{d}+\sum_{i=d+1}^{n}\bm{Y}_{i}^{\top}\bm{Y}_{d}\right)^{4}.$
We consider the binomial expansion term and calculate them separately in Equation 7.18:
$(i)=$ $\mathbb{E}(\sum_{i=d+1}^{n}\bm{Y}_{i}^{\top}\bm{Y}_{d})^{4},\ (ii)=\mathbb{E}\sum_{i=1}^{d-1}\bm{\xi}_{i}^{\top}\bm{Y}_{d}\cdot(\sum_{i=d+1}^{n}\bm{Y}_{i}^{\top}\bm{Y}_{d})^{3},\ (iii)=\mathbb{E}(\sum_{i=1}^{d-1}\bm{\xi}_{i}^{\top}\bm{Y}_{d})^{2}\cdot(\sum_{i=d+1}^{n}\bm{Y}_{i}^{\top}\bm{Y}_{d})^{2},$ (7.19)
$(iv)=$ $\mathbb{E}(\sum_{i=1}^{d-1}\bm{\xi}_{i}^{\top}\bm{Y}_{d})^{3}\cdot\sum_{i=d+1}^{n}\bm{Y}_{i}^{\top}\bm{Y}_{d},\ (v)=\mathbb{E}(\sum_{i=1}^{d-1}\bm{\xi}_{i}^{\top}\bm{Y}_{d})^{4}.$
Since $\mathbb{E}\bm{Y}_{i}=\mathbb{E}\bm{\xi}_{i}=0$, we easily find that Equation 7.19 -(ii)(iv) equal to 0. Next we can get the following equations for Equation 7.19-(iii) after some straightforward calculations.
$\mathbb{E}(\sum_{i=1}^{d-1}\bm{\xi}_{i}^{\top}\bm{Y}_{d})^{2}\cdot(\sum_{i=d+1}^{n}\bm{Y}_{i}^{\top}\bm{Y}_{d})^{2}=$ $\mathbb{E}[\mathbb{E}[(\sum_{i=1}^{d-1}\bm{\xi}_{i}^{\top}\bm{Y}_{d})^{2}\cdot(\sum_{i=d+1}^{n}\bm{Y}_{i}^{\top}\bm{Y}_{d})^{2}\mid\bm{Y}_{d}]]$ (7.20)
$=$ $\mathbb{E}[(d-1)(n-d)(\bm{Y}_{d}^{\top}\mathbf{\Sigma}_{u}\bm{Y}_{d})^{2}]$
$=$ $\mathbb{E}[(d-1)(n-d)((\mathbf{\Sigma}_{u}^{-1/2}\bm{Y}_{d})^{\top}\mathbf{\Sigma}_{u}^{2}(\mathbf{\Sigma}_{u}^{-1/2}\bm{Y}_{d}))^{2}]$
$=$ $(d-1)(n-d)\cdot 2\text{tr }(\mathbf{\Sigma}_{u}^{4})$
$\leq$ $(d-1)(n-d)\cdot O(\text{tr}(\mathbf{\Sigma}_{u}^{2})^{2}).$
By some properties for standard normal random variable, the last inequality holds with some simple calculations shown below.
tr ($\text{tr }(\mathbf{\Sigma}_{u}^{4})$ $=\|\mathbf{\Sigma}_{u}^{2}\|_{F}^{2}=\|\mathbf{\Sigma}_{u}\cdot\mathbf{\Sigma}_{u}\|_{F}^{2}\leq(\|\mathbf{\Sigma}_{u}\|_{F}\cdot\|\mathbf{\Sigma}_{u}\|_{F})^{2}$ (7.21)
$=\|\mathbf{\Sigma}_{u}\|_{F}^{4}=\text{tr }(\mathbf{\Sigma}_{u}^{2})^{2}.$
(ii)If $\bm{X},\bm{Y}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}N(0,\mathbf{I}_{p})$, then
$\mathbb{E}(\bm{X}^{\top}\mathbf{A}\bm{X})^{2}$ $=2\text{tr}(\mathbf{A}^{2})-\text{tr}^{2}(\mathbf{A})\leq\text{tr}(\mathbf{A}^{2}),$ (7.22)
$\mathbb{E}(\bm{X}^{\top}\mathbf{A}\bm{Y})^{4}$ $=\mathbb{E}[\mathbb{E}[(\bm{Y}^{\top}\mathbf{A}\bm{X}\bm{X}^{\top}\mathbf{A}\bm{Y})^{2}\mid\bm{X}]]\leq 2\mathbb{E}[\text{tr}(\mathbf{A}\bm{X}\bm{X}^{\top}\mathbf{A})^{2}]$
$=2\mathbb{E}[(\bm{X}^{\top}\mathbf{A}^{2}\bm{X})^{2}]\leq 4\text{tr}(\mathbf{A}^{4}).$
For Equation 7.19-(1), according to $\sum_{i=d+1}^{n}\bm{Y}_{i}\sim N(0,(n-d)\mathbf{\Sigma}_{u})$ and Equation 7.21-7.22, we have,
$\mathbb{E}(\sum_{i=d+1}^{n}\bm{Y}_{i}^{\top}\bm{Y}_{d})^{4}=$ $\mathbb{E}((\frac{1}{\sqrt{(n-d)}}\mathbf{\Sigma}_{u}^{-1/2}\sum_{i=d+1}^{n}\bm{Y}_{i})^{\top}(\sqrt{n-d}\mathbf{\Sigma}_{u})(\mathbf{\Sigma}_{u}^{-1/2}\bm{Y}_{d}))^{4}$ (7.23)
$\leq$ tr (($\text{tr }((n-d)^{2}\mathbf{\Sigma}_{u}^{4})=(n-d)^{2}O(\text{tr}(\mathbf{\Sigma}_{u}^{2})^{2}).$
Similar with Equation 7.23, for Equation 7.19-(v),
$\mathbb{E}(\sum_{i=1}^{d-1}\bm{\xi}_{i}^{\top}\bm{Y}_{d})^{4}\leq(d-1)^{2}\text{tr }(\mathbf{\Sigma}_{u}^{4})\leq(d-1)^{2}O(\text{tr}(\mathbf{\Sigma}_{u}^{2})^{2}).$ (7.24)
Thus, in combining with the Equation 7.10,
$\quad\mathbb{E}\left(\left(W_{d}-W_{d,0}\right)^{4}\right)$
$=\frac{p^{4}}{2n^{4}(\text{tr }(R^{2}))^{2}}\mathbb{E}\left(\sum_{i=1}^{d-1}\bm{\xi}_{i}^{\top}\bm{Y}_{d}+\sum_{i=d+1}^{n}\bm{Y}_{i}^{\top}\bm{Y}_{d}\right)^{4}$
$\leq\frac{p^{4}}{2n^{4}(\text{tr }(\mathbf{R}^{2}))^{2}}\{(d-1)(n-d)+(n-d)^{2}+(d-1)^{2}\}O(\text{tr}(\mathbf{\Sigma}_{u}^{2})^{2})$
$\leq\frac{p^{4}}{2n^{4}(\text{tr }(\mathbf{R}^{2}))^{2}}n^{2}O(\text{tr}(\mathbf{\Sigma}_{u}^{2})^{2})=O(\frac{1}{n^{2}}).$
By Jensen’s inequality , we get
$\sum_{d=1}^{n}\mathrm{E}\left|W_{d}-W_{d,0}\right|^{3}\leq\sum_{d=1}^{n}\left(\mathrm{E}\left(W_{d}-W_{d,0}\right)^{4}\right)^{3/4}\leq C^{\prime}n^{-1/2},$
for some positive constant $C^{\prime}$, Combining all facts together, we conclude that
$\sum_{d=1}^{n}\left|\mathrm{E}\left\{f\left(W_{d},V_{d}\right)\right\}-\mathrm{E}\left\{f\left(W_{d+1},V_{d+1}\right)\right\}\right|\leq C\beta^{2}n^{-1/2}\log^{3}(n)+C^{\prime}n^{-1/2}\rightarrow 0,$
as $(n,p)\rightarrow\infty$. The conclusion follows.
7.2.6 Proof of Theorem 7(asymptotically independent under $H_{1,local}$)
Proof.  From the proof of Theorem 2 in Feng and Sun (2016), we can find that
$T_{SUM}=\frac{2}{n(n-1)}\sum\sum_{i<j}\bm{U}_{i}^{\top}\bm{U}_{j}+\zeta_{1}^{2}\bm{\theta}^{\top}\mathbf{D}^{-1}\bm{\theta}+o_{p}(\sigma_{n}),$
and according to Lemma 1 with minor modifications, we get the Bahadur representation in $L^{\infty}$ norm,
$n^{1/2}{\mathbf{D}}^{-1/2}(\hat{\bm{\theta}}-\bm{\theta})=n^{-1/2}\zeta_{1}^{-1}\sum_{i=1}^{n}(\bm{U}_{i}+\zeta_{1}{\mathbf{D}}^{-1/2}\bm{\theta})+C_{n}.$
Similar with the proof in Theorem 6, it’s suffice to show the result holds for normal version, i.e. it suffice to show that:
$\|\sqrt{\frac{p}{n}}\sum_{i=1}^{n}\bm{Y}_{i}\|^{2}\text{ and }\|\sqrt{\frac{p}{n}}\sum_{i=1}^{n}(\bm{Y}_{i}+\zeta_{1}{\mathbf{D}}^{-1/2}\bm{\theta})\|_{\infty}^{2},$
are asymptotic independent, where $\{\bm{Y}_{1},\bm{Y}_{2},\cdots,\bm{Y}_{n}\}$ are sample from $N(0,\mathbb{E}\bm{U}_{1}^{\top}\bm{U}_{1})$.
Denote $\sqrt{\frac{p}{n}}\sum_{i=1}^{n}\bm{Y}_{i}:=\bm{\varphi}=(\varphi_{1},\cdots,\varphi_{p})^{\top}$,$\bm{\varphi}_{\mathcal{A}}=(\varphi_{j_{1}},\cdots,\varphi_{j_{d}})^{\top}$, and $\bm{\varphi}_{\mathcal{A}^{c}}=(\varphi_{j_{d+1}},\cdots,\varphi_{j_{p}})^{\top}$, where $\mathcal{A}=\{j_{1},j_{2},\cdots,j_{d}\}$. Then, $S=\|\bm{\varphi}\|^{2}=\|\bm{\varphi}_{\mathcal{A}}\|^{2}+\|\bm{\varphi}_{\mathcal{A}^{c}}\|^{2}$, $M=\|\bm{\varphi}+\sqrt{np}\zeta_{1}{\mathbf{D}}^{-1/2}\bm{\theta}\|_{\infty}=\max_{i\in\mathcal{A}}(\varphi_{i}+\sqrt{np}\zeta_{1}{\mathbf{D}}^{-1/2}\bm{\theta})+\max_{i\in\mathcal{A}^{c}}\varphi_{i}$. From the proof of Theorem 6, we know that $\|\bm{\varphi}_{\mathcal{A}^{c}}\|^{2}$ and $\max_{i\in\mathcal{A}^{c}}\varphi_{i}$ are asymptotically independent. Hence, it suffice to show that $\|\bm{\varphi}_{\mathcal{A}^{c}}\|^{2}$ is asymptotically independent with $\bm{\varphi}_{\mathcal{A}}$.
By Lemma 14, $\bm{\varphi}_{\mathcal{A}^{c}}$ can be decomposed as $\bm{\varphi}_{\mathcal{A}^{c}}=\bm{E}+\bm{F}$, where $\bm{E}=\bm{\varphi}_{\mathcal{A}^{c}}-\mathbf{\Sigma}_{U,\mathcal{A}^{c},\mathcal{A}}\mathbf{\Sigma}^{-1}_{U,\mathcal{A},\mathcal{A}}\bm{\varphi}_{\mathcal{A}},\bm{F}=\mathbf{\Sigma}_{U,\mathcal{A}^{c},\mathcal{A}}\mathbf{\Sigma}^{-1}_{U,\mathcal{A},\mathcal{A}}\bm{\varphi}_{\mathcal{A}}$, $\mathbf{\Sigma}_{U}=p\mathbb{E}\bm{U}_{1}\bm{U}_{1}^{\top}=p\mathbf{\Sigma}_{u}$ , which fulfill the properties $\bm{E}\sim N(0,\mathbf{\Sigma}_{U,\mathcal{A}^{c},\mathcal{A}^{c}}-\mathbf{\Sigma}_{U,\mathcal{A}^{c},\mathcal{A}}\mathbf{\Sigma}^{-1}_{U,\mathcal{A},\mathcal{A}}\mathbf{\Sigma}_{U,\mathcal{A},\mathcal{A}^{c}})$, $\bm{F}\sim N(0,\mathbf{\Sigma}_{U,\mathcal{A}^{c},\mathcal{A}}\mathbf{\Sigma}^{-1}_{U,\mathcal{A},\mathcal{A}}\mathbf{\Sigma}_{U,\mathcal{A},\mathcal{A}^{c}})$ and $\bm{E}$ and $\bm{\varphi}_{\mathcal{A}}$ are independent.
Then, we rewrite
$\|\bm{\varphi}_{\mathcal{A}^{c}}\|^{2}=\bm{E}^{\top}\bm{E}+\bm{F}^{\top}\bm{F}+2\bm{E}^{\top}\bm{F}.$
According the proof of lemma S.7 in Feng et al. (2022a), we have that:
$\mathbb{P}(|\bm{F}^{\top}\bm{F}+2\bm{E}^{\top}\bm{F}|\geq\epsilon\nu_{p})\leq\frac{3}{p^{t}}\rightarrow 0,$
by $d=o(\lambda_{\min}(\mathbf{R})\text{tr}(\mathbf{R}^{2})^{1/2}/(\log p)^{C})$, where $\nu_{p}=[2\text{tr}(\mathbf{R}^{2})]^{1/2}$ , $t=t_{p}:=C\epsilon/8\cdot v_{p}/[\lambda_{\max}(\mathbf{R})\log p]\rightarrow\infty$,$\epsilon_{p}:=(\log p)^{C}/[v_{p}\lambda_{\min}(\mathbf{R})]\rightarrow 0$.
7.3 Some useful lemmas
Lemma 10.
(Lemma A3. in Cheng et al. (2023)) Suppose Assumptions 1-3 holds with $a_{0}(p)\asymp p^{1-\delta}$ for some positive constant $\delta\leq 1/2$. Define a random $p\times p$ matrix $\mathbf{Q}=n^{-1}\sum_{i=1}^{n}R^{-1}_{i}\bm{U}_{i}\bm{U}_{i}^{\top}$ and let $\mathbf{Q}_{jl}$ be the $(j,l)$th element of $\mathbf{Q}$. Then,
(i)$|\mathbf{Q}_{jl}|\lesssim\zeta_{1}p^{-1}|\sigma_{jl}|+O_{p}(\zeta_{1}n^{-1/2}p^{-1}+\zeta_{1}p^{-7/6}+\zeta_{1}p^{-1-\delta/2})$.
(ii)$\mathbf{Q}_{jl}=\mathbf{Q}_{0,jl}+O_{p}(\zeta_{1}p^{-7/6}+\zeta_{1}p^{-1-\delta/2})$, where $\mathbf{Q}_{0,jl}$ is the $(j,l)$th element of
$\mathbf{Q}_{0}=n^{-1}p^{-1/2}\sum_{i=1}^{n}\nu_{i}^{-1}\{\mathbf{D}^{-1/2}\mathbf{\Gamma}U(\bm{W}_{i})\}\{\mathbf{D}^{-1/2}\mathbf{\Gamma}U(\bm{W}_{i})\}^{\top}.$
In addition, $\mathbf{Q}_{0}$ satisfies
[…]
Lemma 11.
(Lemma A4. in Cheng et al. (2023))Suppose Assumptions 1-3 holds with $a_{0}(p)\asymp p^{1-\delta}$ for some positive constant $\delta\leq 1/2$. Then,
(i) $\mathbb{E}\{(\zeta_{1}^{-1}U_{i,j})^{4}\}\lesssim\bar{M}^{2}$ and $\mathbb{E}\{(\zeta_{1}^{-1}U_{i,j})^{2}\}\gtrsim\underline{m}$ for all $i=1,2,\cdots,n$ and $j=1,2,\cdots,p$.
(ii) $\|\zeta_{1}^{-1}U_{i,j}\|_{\psi_{\alpha}}\lesssim\bar{B}$ for all $i=1,2,\cdots,n$ and $j=1,2,\cdots,p$.
(iii )$\mathbb{E}(U_{i,j}^{2})=p^{-1}+O(p^{-1-\delta/2})$ for $j=1,2,\cdots,p$ and $\mathbb{E}(U_{i,j}U_{i,l})=p^{-1}\sigma_{j,l}+O(p^{-1-\delta/2})$ for $1\leq j\not=l\leq p$.
(iv) if $\log p=o(n^{1/3})$,
$\left|n^{-1/2}\sum_{i=1}^{n}\zeta_{1}^{-1}\bm{U}_{i}\right|_{\infty}=O_{p}\{\log^{1/2}(np)\}\text{ and }\left|n^{-1}\sum_{i=1}^{n}(\zeta_{1}^{-1}\bm{U}_{i})^{2}\right|_{\infty}=O_{p}(1).$
Lemma 12.
(Nazarov’s inequality) Let $\bm{Y}_{0}=(Y_{0,1},Y_{0,2},\cdots,Y_{0,p})^{\top}$ be a centered Gaussian random vector in $\mathbb{R}^{p}$ and $\mathbb{E}(Y_{0,j}^{2})\geq b$ for all $j=1,2,\cdots,p$ and some constant $b>0$, then for every $y\in\mathbb{R}^{p}$ and $a>0$,
$\mathbb{P}(\bm{Y}_{0}\leq y+a)-\mathbb{P}(\bm{Y}_{0}\leq y)\lesssim a\log^{1/2}(p).$
Lemma 13.
(Theorem 2 in Chernozhukov et al. (2015)) Let $\bm{X}=\left(X_{1},\ldots,X_{p}\right)^{\top}$ and $\bm{Y}=\left(Y_{1},\ldots,Y_{p}\right)^{\top}$ be centered Gaussian random vectors in $\mathbb{R}^{p}$ with covariance matrices $\mathbf{\Sigma}^{X}=$ $\left(\sigma_{jk}^{X}\right)_{1\leq j,k\leq p}$ and $\mathbf{\Sigma}^{Y}=\left(\sigma_{jk}^{Y}\right)_{1\leq j,k\leq p}$, respectively. In terms of $p$,
$\Delta:=\max_{1\leq j,k\leq p}\left|\sigma_{jk}^{X}-\sigma_{jk}^{Y}\right|,\text{ and }a_{p}:=\mathrm{E}\left[\max_{1\leq j\leq p}\left(Y_{j}/\sigma_{jj}^{Y}\right)\right].$
Suppose that $p\geq 2$ and $\sigma_{jj}^{Y}>0$ for all $1\leq j\leq p$. Then
$\sup_{x\in\mathbb{R}}\left|\mathrm{P}\left(\max_{1\leq j\leq p}X_{j}\leq x\right)-\mathrm{P}\left(\max_{1\leq j\leq p}Y_{j}\leq x\right)\right|$
$\quad\leq C\Delta^{1/3}\left\{\left(1\vee a_{p}^{2}\vee\log(1/\Delta)\right\}^{1/3}\log^{1/3}p,\right.$
where $C>0$ depends only on $\min_{1\leq j\leq p}\sigma_{jj}^{Y}$ and $\max_{1\leq j\leq p}\sigma_{jj}^{Y}$ (the right side is understood to be 0 when $\Delta=0$ ). Moreover, in the worst case, $a_{p}\leq\sqrt{2\log p}$, so that
$\sup_{x\in\mathbb{R}}\left|\mathrm{P}\left(\max_{1\leq j\leq p}X_{j}\leq x\right)-\mathrm{P}\left(\max_{1\leq j\leq p}Y_{j}\leq x\right)\right|\leq C^{\prime}\Delta^{1/3}\{1\vee\log(p/\Delta)\}^{2/3},$
where as before $C^{\prime}>0$ depends only on $\min_{1\leq j\leq p}\sigma_{jj}^{Y}$ and $\max_{1\leq j\leq p}\sigma_{jj}^{Y}$.
Lemma 14.
(Theorem 1.2.11 in Muirhead (2009))Let $\bm{X}\sim N(\bm{\mu},\mathbf{\Sigma})$ with invertible $\mathbf{\Sigma}$, and partition $\bm{X},\bm{\mu}$ and $\mathbf{\Sigma}$ as
$\bm{X}=\left(\begin{array}[]{c}\bm{X}_{1}\\ \bm{X}_{2}\end{array}\right),\quad\bm{\mu}=\left(\begin{array}[]{l}\bm{\mu}_{1}\\ \bm{\mu}_{2}\end{array}\right),\quad\mathbf{\Sigma}=\left(\begin{array}[]{cc}\mathbf{\Sigma}_{11}&\mathbf{\Sigma}_{12}\\ \mathbf{\Sigma}_{21}&\mathbf{\Sigma}_{22}\end{array}\right).$
Then $\bm{X}_{2}-\mathbf{\Sigma}_{21}\mathbf{\Sigma}_{11}^{-1}\bm{X}_{1}\sim N\left(\bm{\mu}_{2}-\mathbf{\Sigma}_{21}\mathbf{\Sigma}_{11}^{-1}\bm{\mu}_{1},\mathbf{\Sigma}_{22\cdot 1}\right)$ and is independent of $\bm{X}_{1}$, where $\mathbf{\Sigma}_{22\cdot 1}=$ $\mathbf{\Sigma}_{22}-\mathbf{\Sigma}_{21}\mathbf{\Sigma}_{11}^{-1}\mathbf{\Sigma}_{12}$.